### v1

# Query Rewriting via LLMs

Abstract
Query rewriting is a classical technique for transforming complex declarative SQL queries into “lean” equivalents that are conducive to (a) faster execution from a performance perspective, and (b) better understanding from a developer perspective. The rewriting is typically achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system. In recent times, LLM-based techniques have also been mooted, but they are prone to both semantic and syntactic errors.
We investigate here, how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency. Our study shows that these goals can be progressively achieved through incorporation of (a) an ensemble suite of basic prompts, (b) database-sensitive prompts via redundancy removal and selectivity-based rewriting rules, and (c) LLM token probability-guided rewrite paths. Further, a suite of statistical and logic-based tools can be used to guard against errors produced by the model.
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from multiple benchmarks on contemporary database platforms. The results show significant improvements over SOTA rewriting techniques – for instance, on TPC-DS, LITHE constructed productive ($>$1.5x speedup) rewrites for two-thirds of the query suite, delivering four times more coverage than SOTA. Further, the geometric mean of its estimated execution speedups was an order-of-magnitude jump over SOTA performance. In essence, LITHE offers a potent and robust LLM-based intermediary between enterprise applications and database engines.
1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools. A compelling case in point is the blog-processing query [4] shown in Figure 2 – this query was produced by the popular Entity Framework [21] mapper. It computes a daily summary of rating metrics for highly-rated blogs. This seemingly complex nested query can be equivalently rewritten in the “lean” flat version shown in Figure 2.
SELECT t.Key, sum(t.Rating) AS PostRating,
(SELECT sum(b0.Rating)
FROM (SELECT p0.PostId, p0.BlogId, p0.Content,
p0.CreatedDate, p0.Rating, p0.Title,
b1.BlogId AS BlogId0, b1.Rating AS Rating0,
b1.Url, p0.day AS Key
FROM Posts AS p0 INNER JOIN Blogs AS b1 ON
p0.BlogId = b1.BlogId
WHERE b1.Rating > 5) AS t0
INNER JOIN Blogs AS b0 ON t0.BlogId = b0.BlogId
WHERE t.Key = t0.Key ) AS BlogRating
FROM (SELECT p.Rating, p.day AS Key
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5) AS t
GROUP BY t.Key;
Figure 1: Complex SQL Representation
SELECT p.day AS Key, SUM(p.Rating) AS PostRating,
SUM(b.Rating) AS BlogRating
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5
GROUP BY p.day;
Figure 2: Lean Equivalent Query
Albeit less common, we observe such superfluity even in the hand-crafted expert query domain – for instance, TPC-DS Query Q23 includes a sub-query that identifies customers with the maximum store sales in a four-year period; this sub-query has a redundant join of store_sales with customer despite the FK-PK relationship between these tables, as highlighted in [25].
Transforming bloated queries to lean equivalents in the application source code is beneficial from multiple perspectives: First, rewrites can improve code readability, a major concern in industrial contexts. Second, query optimizers are expected to, in principle, remove bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor query performance. In fact, both examples discussed above were implemented as such by the current PostgreSQL optimizer (v16) [1], without any change. Therefore, a query rewriter can serve as an effective and non-invasive mechanism for delivering good performance despite inherent optimizer limitations.
Given this motivational context, a viable SQL-to-SQL query transformer should satisfy the following criteria: (1) The transformed query should be semantically equivalent to the original; (2) There should ideally be performance improvement due to the rewrite, but at the very least, no regression in query performance; and (3) The transformation overheads must be practical for deployment.
Prior Work
A variety of innovative Rule-based (e.g. Learned Rewrite [38]) and Model-based (e.g. Gen-Rewrite [19]) approaches have been proposed in the literature for SQL query transformations. However, as explained later in the related work section (Section 8), the benefits of these state-of-the-art (SOTA) techniques are curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space.
An alternative and extremely simple strategy would be to take advantage of the large body of recent literature on LLM-based Text-to-SQL translations (e.g. [17, 35, 33]). That is, first invoke an LLM to transform the original SQL to text, and then employ a Text-to-SQL tool to produce quality SQL. However, as discussed later in Section 7, we found this approach to be highly prone to semantic errors and lacking execution benefits. This is perhaps to be expected given the focus of Text-to-SQL tools on query formulation, orthogonal to our objective of performance enhancement. Therefore, the SQL-to-SQL problem merits a fresh assessment from first principles.
The LITHE Rewriter
In this study, we carry out a calibrated investigation of how LLMs could be reliably used to produce performant rewrites of complex SQL queries. The proposed techniques, described below, have been implemented and evaluated in the LITHE (LLM Infused Transformations of Hefty queries) rewriting system developed by our group.
Basic Prompts. We begin with an ensemble suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs. Interestingly, we find that more information is not necessarily better wrt rewriting quality, and the best prompt granularity is query-specific. Moreover, this simple prompt ensemble was found to deliver performance similar to the SOTA techniques.
Database-sensitive Prompts. To help the model adapt to different query patterns and structures, we next introduce rules in the prompts. But our choice and usage of rules are markedly different from those used in prior work. First, our rules are invoked directly in query space. The LLM therefore gains the latitude to generalize the rule usage to a wide range of queries. This is in contrast to the hardwired and narrow rule application mechanisms (e.g. Calcite [8] rules) typically used in existing rewrite systems, which operate in plan space. Second, to minimize rule application overheads, we incorporate only a handful of rules that serve to eliminate the primary causes of poor query performance. Whereas current rewrite systems employ a large number of rules in their processing. In particular, we work with two classes of rules:
1. Redundancy Removal Rules: These rules are meant to eliminate repeated and redundant computations of the same output.
2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
Token Probability Driven Rewrites. The above rewrite options are restricted to the standard prompt interface. Additionally, we leverage the rich telemetry provided by LLMs – in particular, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation. MCTS has previously demonstrated significant effectiveness in code generation exercises [36].
Experimental Evaluation
Our experiments are carried out on a variety of synthetic and real benchmarks: TPC-DS [11], DSB [12], JOB [15], StackOverflow [20] and Archer [37], covering a broad spectrum of databases and queries. Further, anonymized versions, where the table and column names are altered to not convey any semantic meaning about their contents, are also evaluated to assess robustness to databases not seen by the LLM during its training phase. These databases are hosted on PostgreSQL and GPT-4o is used as the LLM.
We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [38], LLM-R2 [18], GenRewrite [19]) as well as a baseline LLM prompt [19]. The primary metrics are the execution cost reductions and the rewriting overheads. For LLM-based techniques, the number of tokens in the prompts is also monitored since the financial charges for LLM usage are typically dependent on this number. To understand the impact of the various components of LITHE, a systematic ablation study is carried out. Finally, the performance on alternative database and LLM platforms is also evaluated.
Verifying Semantic Equivalence. The general problem of proving query equivalence has long been known to be computationally hard [6]. However, for restricted classes of queries, various logic-based equivalence-proving tools are available (e.g. Cosette [10], SQL-Solver [13], and QED [30]). In our LITHE system, we use QED [30] to verify rewrite correctness for queries within its scope. The remainder are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases using small sampled versions of the original database. We also provide the user with an option to test result equivalence using the original database.
Results. Our experiments demonstrate that LITHE achieves semantically correct transformations that deliver a significant reduction in estimated execution times for many queries. Moreover, LITHE performed better than or at least as well as SOTA on all the benchmarks. As a compelling case in point, LITHE constructed “productive” (¿1.5x speedup) rewrites for about two-thirds of the complete TPC-DS query suite, delivering four times more coverage than SOTA techniques. Further, the GM (Geometric Mean) of the reduction in query execution costs was 30.6, an order-of-magnitude jump over that offered by SOTA.
In essence, LITHE offers a potent LLM-based intermediary between enterprise applications and database engines.
Contributions
In summary, our study makes the following contributions:
  1. 1.
    Assesses LLM suitability for SQL-to-SQL transformation.
  2. 2.
    Transforms directly in query space instead of a plan space intermediate representation, leading to robust rewrites.
  3. 3.
    Incorporates potent database-sensitive rules in LLM prompts covering schematic and statistical dimensions.
  4. 4.
    Leverages LLM token probabilities to guide navigation of the rewrite search space and minimize LLM errors.
  5. 5.
    Evaluates rewriting robustness over a broad range of database environments and benchmarks.
  6. 6.
    Demonstrates substantive performance benefits, well beyond SOTA, from LLM-powered query rewrites.
2 LITHE Overview
The LITHE architecture, illustrated in Figure 3, consists of two main components: A prompt-based rewriting pipeline, and a token probability-driven rewrite exploration.
Prompt-based rewriting pipeline. This pipeline comprises four modules: (1) LLM Prompting, (2) Syntax Verification, (3) Query Costing, and (4) Semantic Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output is checked by the syntax verification component – if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively until a valid SQL query is obtained. Then, the execution cost of this version is evaluated – if more than the cost of the original query, the rewrite is discarded. Otherwise, the rewrite is tested for semantic equivalence. If a match is established, the rewrite is returned; otherwise, the original query itself is returned. Finally, to prevent runaway situations, a threshold is set on the number of syntax correction attempts (5 in our experiments) and the original query is returned if this threshold is breached.
The above process is repeated for each of the different prompts described in Sections 3 and 4. Then, the prompt providing the least-cost rewrite is returned. Since each prompt is evaluated within a fresh LLM context, there is no dependency on the order in which they are assessed.
Within the pipeline, LITHE uses (a) the query parser from the database engine for syntax verification, (b) the query optimizer for cost estimation, and (c) the techniques described in Section 6.3 to check semantic equivalence.
Figure 3: High-level architecture of LITHE
Token probability-driven rewrite. The pipeline-chosen prompt is used as an input to a Monte-Carlo tree search (MCTS)-based procedure to further improve the query rewrite quality. Specifically, multiple exploration paths are followed whenever the LLM lacks high confidence in the predicted token (details in Section 5).
2.1 Performance Framework
As mentioned in the Introduction, the majority of our experiments are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine.
We define a productive rewrite (PR) as one that improves the query’s performance by at least 1.5 times wrt the execution times estimated by the database engine’s query optimizer – this aggressive choice of threshold is to make the rewrite outcome desirable from an industrial perspective. Given a query workload, the set of feasible productive rewrites (FPR) consists of the queries for which PR are known to be feasible by some technique or the other. Note that it may be inherently impossible to produce PR on queries that are structurally simple or efficiently written by human experts. Finally, a maximum productive rewrite (MPR) is when the PR speedup of the rewrite matches the best known PR on this query.
The overall benefit provided by a rewriting tool is quantified by the number of PR obtained on the query workload. Additionally, we also measure SpeedupGM, the geometric mean (GM) of the speedups obtained on the FPR queries in this workload. (The geometric mean is used to ensure equal weight for all queries in the workload, irrespective of their relative durations.)
2.2 Query Micro-benchmark
To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft productive rewrites. These human rewrites deliver a SpeedupGM of 13.9 on the micro-benchmark, serving as an informal “upper-bound” on what is computationally attainable. Later, in Section 7, we extend the evaluation to a suite of complete benchmarks.
3 Basic Prompts
In this section, we explore the simplest interface to LLMs, namely prompting, for query rewriting. We evaluate four basic prompts, illustrated in Figure 4, which cover a progressive range of detail in the instructions and test the effectiveness of the LLM’s base knowledge.
Figure 4: Templates used for Basic Prompts
Prompt 1: This is the baseline prompt used in [19], which simply asks the LLM to rewrite a given query to improve performance.
Prompt 2: Explicit instructions are included to maintain semantic and functional equivalence while rewriting.
Prompt 3: Verbose instructions are given to rewrite the query, providing step-by-step guidance to the LLM to think rationally. It is first asked to pick out potential inefficiencies in the input query, and then tasked to identify approaches to address these inefficiencies. Finally, it is instructed to apply the identified solution. Essentially, the prompt tries to make the LLM reason akin to a human expert.
Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once. The idea is to break down the complex instructions given in Prompt 3 into digestible steps that help the LLM focus on individual tasks.
Performance
The performance of the four prompt templates on the micro-benchmark (where all 10 queries are in FPR) is shown in Table 1. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that each prompt does the best on certain queries – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. While this ensemble approach (Row 5 in Table 1) raises the PR to 6, there remain four queries that could not be productively rewritten by the prompts alone.
Table 1: Performance of Basic Prompts
Prompt # PR SpeedupGM
Prompt 1 3 2.7
Prompt 2 4 1.57
Prompt 3 3 1.37
Prompt 4 2 1.27
Prompt Ensemble 6 3.6
SOTA Ensemble 5 3.4
The SpeedupGM, shown in the last column, for each individual prompt is less than 3, while that of the ensemble reaches 3.6. But these speedups, while productive, are all lower than those of the human rewrites, that is, they are not MPR.
Finally, as another baseline, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform. They delivered (Row 6 in Table 1) 5 productive rewrites with a GM of 3.4, indicating the wide gap between the current state-of-the-art and what is possible.
4 Database-Sensitive Prompts
As discussed above, basic prompting needs to be improved on two fronts: (1) Ensuring productive rewrites where feasible; and (2) Maximizing the impact of these productive rewrites. To address these issues, we incorporate database domain knowledge. Specifically, we design a one shot-based prompting template, augmented with a set of database-aware rewrite rules. The rules are based on common practices followed by DBAs that are widely applicable, and augmented with precise instructions and useful examples to help guide the LLM in the rewriting process.
As a proof of concept, we explore two categories of rewrites here: (a) Rules that eliminate redundancy in the input queries; and (b) predicate selectivity-based rules that implicitly guide, via query space reformulations, the query optimizer towards efficient query execution plans. Of course, this basic set of rules can be expanded further, but as shown by our experiments, even this minimal set is capable of delivering substantive improvements over a broad set of database environments.
4.1 Redundancy Removal
There are different types of redundancy that can occur in a SQL query – repeated computations, superfluous filter predicates, unnecessary joins, etc. Rules R1 through R4 in Table 2 are designed to tackle such redundancies. For instance, the redundant PK-FK join highlighted in the Introduction is addressed by rule R4 in Table 2 (the relevant schematic information is also provided in the prompt).
Table 2: Rules for Database-sensitive prompts
Redundancy Removal Rules
R1 Use CTEs (Common Table Expressions) to avoid repeated computation.
R2 When multiple subqueries use the same base table, rewrite to scan the base table only once.
R3 Remove redundant conjunctive filter predicates.
R4 Remove redundant key (PK-FK) joins.
Statistics-based Rules
R5 Choose EXIST or IN from subquery selectivity (high/low).
R6 Pre-filter fact tables in a CTE using dimension tables with low selectivities. Retain dimension table filters in main query. Do not create explicit join statements.
The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are enumerated in the Appendix A3. Note that this prompt template allows for only a single rule to be present in the prompt. This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
Figure 5: Templates for Database-sensitive Prompts
Performance
The performance improvement achieved on the micro-benchmark by an ensemble that adds the redundancy-removing prompts to the basic set (Section 3) is shown in Table 3. We observe that the PR increases to 8, and SpeedupGM grows to 4.8.
Table 3: Performance with Redundancy Removal Rules
Prompt # PR SpeedupGM
[…]
A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on. This approach would benefit queries with multiple types of redundancies. However, it also increases rewrite overheads due to repeated prompt applications. So, for simplicity, we have chosen to process them individually rather than cumulatively.
Figure 6: Example Queries illustrating Rule 5
4.2 Selectivity-based Guidance
We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. Based on such arguments, rules R5 and R6, which pre-filters fact tables for dimension tables with low selectivity values, are included in Table 2. Note that, a specific instruction to not create explicit joins had to be added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
The input prompt for these rules, as shown in Figure 5(b), is modified to include the following:
  1. 1.
    Selectivities of columns appearing in WHERE and JOIN clauses, as obtained from the database engine.
  2. 2.
    Clause rewrite rules and instructions based on statistics.
  3. 3.
    Examples relevant to the chosen rewrite rules.
Performance
The performance improvements following addition of selectivity-guided prompts are shown in Table 4. We observe that PR are now obtained for all 10 micro-benchmark queries. Moreover, the resulting SpeedupGM increases to 11.8, quite close to the human upper bound of 13.9.
Table 4: Performance of Metadata-infused Prompts
Prompt # PR SpeedupGM
[…]
A key point to note in closing is that rules R1 through R6 not only bring additional queries into the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule, and this option is examined in Section 7.3.3.
5 Token Probability Driven Rewrite
A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the potential output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs. In order to have a robust approach for such cases, we take inspiration from the code generation literature [36]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [24]. The root node of the tree corresponds to the initial prompt. An edge from a parent node to a child represents a possible token generated by the LLM and is associated with a value denoting the probability of generating this token given the path taken thus far. Here, each edge can be considered as an action of the MDP. A node is considered terminal if the incoming edge corresponds to the “;” token, signalling end of the textual query.
The state of a node $n$ is represented by the partial rewrite created by following the path from the root to $n$ – it is obtained by concatenating the tokens on this path. The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query.
Given that the vocabulary sizes of LLMs are upward of hundreds of thousands of tokens, it may become very expensive (in both financial and computational costs) to construct the entire tree. It is therefore essential to significantly reduce the token search space while exploring the tree for valid rewrites. This is precisely the purpose of MCTS which applies an Upper Confidence Bound (UCB) heuristic [26] to identify the best paths in a tree without computing the entire tree.
5.1 MCTS Search Process
The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
1. Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure). Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1). Actions are picked using a UCB that balances exploration and exploitation. The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
Specifically, given a node $n$ and a set of possible actions $a\in A$, the next node in this traversal is chosen as:
$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
where $UCB$ is a heuristic sourced from [26], modified to reflect our formulation where values are associated with nodes rather than edges in the tree. It is defined as follows:
$UCB(n,a)=V(n^{\prime})+\beta(n)*P_{LLM}(a|n.state)*\frac{\sqrt{log(visits[n])}}{1+visits[n^{\prime}]}$ (2)
Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3). The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often. Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process. $\beta$ is a function that controls the balance between exploration and exploitation. It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore. $\beta$ is defined as:
$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
Algorithm 1 Token-augmented Rewrite
1:$root$         # Start State
2:$k$            # Maximum number of child node expansions
3:$\theta$            # Probability threshold for node expansion
4:$iter_{max}$     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
7:     $visits[root]$leftarrow $visits[root]+1$
8:     $n_{cur}$leftarrow $root$
9:     # Stage 1: Selection
10:     while $len(n_{cur}.children)>0$ do
11:         $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$
12:         $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
13:     # Stage 2: Expansion
14:     $expand\leftarrow$ True
15:     while $expand$ and ‘;’ $\notin n_{cur}.state$ do
16:         $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
17:         if $P_{next}[0]\leq\theta$ then
18:              for $token\in tokens_{next}$ do
19:                  $n_{new}\leftarrow$ new $Node$ with State $n_{cur}.state\cdot token$
20:                  Append $n_{new}$ to $n_{cur}.children$
21:              $expand\leftarrow$ False
22:         else
23:              $n_{cur}.state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$
24:     # Stage 3: Simulation - Expand from $n_{cur}$ to full rewrite
25:     $query\leftarrow$ GreedyExpand($n_{cur}$)
26:     $v\leftarrow$ ComputePotential($query$)
27:     $Potential[query]\leftarrow v$
28:     # Stage 4: Backpropagation
29:     while $n_{cur}\neq\text{Null}$ do
30:         $V[n_{cur}]\leftarrow\max(V[n_{cur}],v)$
31:         $n_{cur}\leftarrow\text{Parent}(n_{cur})$
32:# Return valid rewrite with maximum $Potential>1$
33:if $\exists q\in Potential\mid Potential[q]>1$ then
34:     return $q$ having maximum value of $Potential[q]$
35:else
36:     return the original query
2. Expansion: The second stage is used to expand the unprocessed node $n_{cur}$ chosen by the Selection stage. Specifically, it retrieves from the LLM the top $k$ probable next tokens from $n_{cur}$’s state (Line 12), and expands the decision tree by adding $k$ new child nodes corresponding to these tokens. To make the expansion tractable, multiple child nodes are added only if the probability of the highest token falls below a threshold $\theta$ (Line 13). In other words, when the highest token probability is below $\theta$, it means that the LLM itself is unsure of what the next token should be and therefore it is worth exploring additional options. On the other hand, if the highest token probability is greater than $\theta$, then the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e., completes a query rewrite).
3. Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$. This node is expanded in a greedy fashion, based on the highest-probability tokens until a terminal node is reached (Line 21). Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23). For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides with respect to the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value.
After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is backpropagated to all its ancestor nodes. An ancestor node’s $V$ value is updated if and only if the new value is higher than the existing value.
Rewritten Query. At the end of all iterations, $q\in Potential$ with highest value $Potential[q]$ that is greater than 1 is returned as the rewritten query (Lines 29–30). In case no such rewrite exists, implying that all the valid rewrites are slower than the original query, the original query itself is returned (Line 32).
5.2 Input Prompt to MCTS
The root state in Algorithm 1 corresponds to the state just after the prompt is fed to the LLM. One way to use this algorithm is to execute it for all the various prompts discussed in the previous sections, and choose the rewrite that provides the best performance. This, however, is expensive both from the aspect of query rewrite time, as well as the number of LLM tokens used. To minimize these costs, the LITHE workflow first selects, given a query, the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4. It then employs this prompt to initiate the MCTS-based rewrite. In case no prompt provides a lower-cost rewrite, baseline Prompt 1 of Section 3 is used as the fallback option.
Later, in Section 7.3.3, we show how the overheads incurred can be further reduced via a classifier construction.
5.3 Performance
The performance results with MCTS-based rewrites are shown in Table 5 for the micro-benchmark. We observe gratifyingly that the SpeedupGM now attains the human speedup of 13.9.
Table 5: Performance of MCTS-based rewrites
# PR SpeedupGM
LITHE 10 13.9
6 Implementation Choices
In this section, we briefly discuss the design choices made in our implementation of LITHE.
6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response, with a higher temperature resulting in the outputs being more varied and less predictable. While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible. Hence we set the temperature to 0 which causes the model to do greedy sampling to select the next token.
The hyperparameters used by LITHE for MCTS are as follows: The maximum number of iterations $iter_{max}$ is set to 8, expansion threshold $\theta$ is 0.7, and number of expansions $k$ is 2. The values of $c_{base}$ and $c$ were set to 10 and 4, respectively. These settings were determined after an empirical evaluation of the various trade-offs, providing a robust balance between efficiency and quality.
Finally, we run the prompt pipeline for a maximum of 5 iterations to try and fix any syntax errors created during the rewrite process.
6.2 Query Efficiency Characterization
The LITHE workflow requires computing the performance of several candidate query rewrites. In this scenario, explicitly executing each rewrite on the database to evaluate its runtime benefit may become impractically slow. This is especially true for queries where even a single execution is time-consuming. We therefore leverage the query cost, as estimated by the optimizer, to evaluate the performance of query rewrites. It is common knowledge that these estimates are often inaccurate [16] – however, they are expected to be generally reflective of the actual run times, especially on modern industrial-strength database systems. Moreover, our comparison of costs is only in a relative sense – between the original query and its rewrite, so errors in the absolute values are tolerable as long as similar errors are made in all estimates. These costs can therefore serve as a computationally cheap and meaningful measure to distinguish performance among rewrites.
6.3 Query Equivalence Testing
We use a multi-stage approach, described below, to test semantic equivalence between the original query and a candidate rewrite.
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g. Cosette[10], SQL-Solver [13], QED [30]). In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [34] for maintaining join integrity in the sample, (2) addition of synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Appendix A2.
3. Result Equivalence on the Entire Database. Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself. However, as mentioned above, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
7 Experimental Evaluation
In this section, we report on LITHE’s performance profile. We first describe the experimental setup, including comparative baselines, query suites and evaluation platforms. Then we present the speedup results for both aggregate benchmark and individual queries, followed by characterization of the rewrite overheads in computational and financial terms. We finally discuss the impact of alternative platforms wrt database schema, database engine and LLMs.
Rewrite Baselines. We compare LITHE with a collection of contemporary rewrite techniques, collectively referred to as SOTA – the details of these techniques are provided in Section 8. Specifically, the SOTA collection consists of the following approaches:
  1. 1.
    Baseline LLM prompt [19]: This is Prompt 1 from Section 3.
  2. 2.
    Learned Rewrite [38], a purely rule-based rewriter.
  3. 3.
    LLM-R2 [18], an LLM-guided rule-based rewriter.
  4. 4.
    GenRewrite [19], a purely LLM-based rewriter.
Given an input query, each of the SOTA approaches is independently invoked to perform a rewrite, and the rewrite with the best performance is used as the baseline for comparison. Note that these approaches may generate rewrites that regress the performance [19]. We discard such backsliding rewrites using, similar to LITHE, the Query-Costing module of the database engine.
Query Set. Our evaluation uses the FPR subsets from the following diverse set of benchmarks, amounting to 100+ queries:
  1. 1.
    TPC-DS [11]: A synthetic OLAP benchmark modeling retail supplier environment with complex analytical queries.
  2. 2.
    DSB [12]: A TPC-DS variant with complex data distributions and additional query templates featuring many-to-many joins and non-equi-joins.
  3. 3.
    ARCHER [37]: A Text-to-SQL benchmark spanning 10 databases with availability of ground-truth SQL queries.
  4. 4.
    JOB [16]: An optimizer stress-test benchmark featuring queries with large and complex join graphs.
  5. 5.
    StackOverflow [20]: A real-world benchmark with query templates modeling questions and answers from experts. A random instance of each template is taken.
Testbed. The majority of our experiments are carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA. Variations on this platform are considered in Section 7.4.
Metrics. For each benchmark and each rewrite technique, we identified the number of queries for which a PR (productive rewrite) could be constructed. Subsequently, we computed the GM of the speedups obtained by each technique over the set of FPR (feasible productive rewrites). Finally, recall that if a rewrite achieves the maximum known speedup for the query, it is referred to as an MPR (Maximum Productive Rewrite).
From the investment perspective, we measured the average rewrite time per query, and for the LLM-based techniques, the number of tokens used in the rewrite process.
7.1 Rewrite Quality
Table 6: Comparing LITHE with SOTA on FPR queries
Benchmark (FPR) # PR SpeedupGM
LITHE SOTA LITHE SOTA
TPC-DS (65) 65 16 30.6 2.1
DSB (20) 20 3 22.6 1.3
ARCHER (25) 25 19 2.2 1.8
JOB (5) 5 2 1.8 1.3
StackOverflow (2) 2 1 9.0 7.5
7.1.1 Benchmarks
The speedup delivered by LITHE and SOTA on the FPR queries of the various benchmarks are shown in Table 6. The good news is that for TPC-DS and its DSB variant, LITHE produces PR for all the FPR queries – 65 in TPC-DS and 20 in DSB. In fact, most of these rewrites are highly productive as the SpeedupGM exceeds 30 for TPC-DS and 20 for DSB! Further, the PR coverage is far superior in comparison to SOTA– by a factor of 4 (TPC-DS) and 6 (DSB). Finally, the SpeedupGM of LITHE is higher by more than an order-of-magnitude wrt SOTA.
Figure 7: Rewritten TPC-DS Q88 produced by LITHE
7.1.2 Query Readability
We now turn our attention to assessing how the readability of the queries is impacted by rewrites. That is, we wish to test whether the common wisdom that “efficient queries are more readable” holds for our rewrites. For this purpose, the number of joins in the query is used as a rudimentary metric to quantify query readability. The intuition is that a query with more joins is harder for the user to parse and understand.
Table 7 presents the average number of joins in the original and rewritten queries for both SOTA and LITHE. The results clearly demonstrate that LITHE significantly reduces the number of joins, especially for TPC-DS, DSB and ARCHER benchmark, thereby enhancing query readability. Notably, for the DSB benchmark, SOTA actually worsened the readability, while LITHE maintained superior readability compared to the original query. However, for the JOB and StackOverflow benchmarks, where optimization opportunities were limited, the readability remained unchanged. Importantly, LITHE never performs worse than the original or SOTA-rewritten queries in terms of readability, always maintaining or improving the metric.
Figure 7 shows the rewritten query for Q88 in which the number of joins reduced from 24 to 3. Overall, LITHE proves to be a reliable and effective tool for query rewriting, delivering both better performance and improved readability.
Table 7: LITHE vs SOTA readability comparison
Benchmark (FPR) Average # JOINS
LITHE SOTA Original
TPC-DS (65) 4.1 4.4 4.4
DSB (20) 4.8 5.3 4.95
ARCHER (25) 0.8 1 1.44
JOB (5) 7 7 7
StackOverflow (2) 8 8 8
Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of FPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting. Nevertheless, LITHE continues to achieve PR in all of this limited set, whereas SOTA misses quite a few opportunities. Further, the SpeedupGM of LITHE is visibly better than SOTA.
Figure 8: Speedups achieved by LITHE and SOTA
Figure 9: Rewritten TPC-DS Q90
Figure 10: Query execution times on TPC-DS 100GB
7.1.3 Individual Queries
The above results were for entire benchmarks. We now drill down into the performance at the granularity of individual queries. Due to space limitations, we focus only on the TPC-DS benchmark here.
Figure 8 compares the performance of LITHE and SOTA on each of the 65 FPR queries in TPC-DS – note that the speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale. We first observe that, gratifyingly, rewrites are indeed capable of providing dramatic speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
As another example, Figure 10 shows a rewrite where only LITHE achieves 2X speedup. Here, the two subqueries in the original query (each with 3 joins) are replaced by a common set of joins, and the aggregations are conditionally computed using CASE statements.
Second, even among the 16 PR queries of SOTA, there are instances (e.g. Q6, Q47) where LITHE delivers substantially faster speedups than SOTA. Overall, across these 16 queries, LITHE attains a GM of 28.7 as compared to SOTA’s GM of 19.5.
Finally, note that LITHE matched or improved on SOTA for all the FPR queries.
7.1.4 Query Execution Times.
Thus far, we had considered optimizer estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 10 shows this metric (on a $\mathbf{log_{10}}$ scale) for the original query, LITHE, and SOTA on a representative set of queries executed on the TPC-DS 100GB database.
We observe that substantial benefits are achieved, even exceeding order-of-magnitude improvements in some cases. Of course, as expected, these speedups may not directly match the optimizer-estimates because of the inherent imprecision in optimizer models. For example, the 5 orders-of-magnitude speedup estimated for Q41 decreases to a still respectable 3 orders-of-magnitude. The opposite also occurs: the 2 orders-of-magnitude speedup estimated for Q45 now becomes a huge 4 orders-of-magnitude.
7.2 Ablation Analysis
7.2.1 Components of LITHE
A natural question at this stage is the role of the various techniques in LITHE towards achieving its large performance benefits. This analysis is captured in Table 8 which lists the PR and MPR for each technique when invoked in isolation. The table also shows the cumulative number of PR and MPR when the different techniques are combined in the order of listing.
Table 8: Ablation Analysis
Isolated Cumulative
# PR # MPR # PR # MPR
Basic Prompts (Section 3) 19 13 19 13
Rules R1 — R4 (Section 4.1) 17 12 23 15
Rules R5, R6 (Section 4.2) 44 34 60 48
MCTS (Section 5) 65 65 65 65
We observe in the table that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) are capable of achieving maximum speedups for around a quarter (15/65) of the FPR query set. When the statistics-infused Rules (R5, R6) are added, this fraction jumps to about three-quarters (48/65). And finally, when MCTS is included, we reach full coverage of the FPR.
Table 9: Rewrite Overheads of LITHE and SOTA on FPR queries
Benchmark Avg. Time (min) Avg. Tokens Avg. Cost USD
LITHE SOTA LITHE SOTA LITHE SOTA
TPC-DS 4.6 1.7 22390 20076 0.055 0.050
DSB 8.7 4.0 22203 15699 0.055 0.039
ARCHER 2.3 0.6 7284 5465 0.018 0.013
JOB 6.1 1.3 22958 13692 0.057 0.034
StackOverflow 7.1 3.2 22634 12759 0.056 0.032
7.2.2 Database-Sensitive Prompts.
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads. When this experiment was conducted, the SpeedupGM dropped precipitously to a paltry 2.3, a far cry from the 30.6 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
7.3 Rewrite Overheads
Having established the ability of LITHE to deliver performance-beneficial rewrites, we now turn our attention to the overheads incurred in this rewriting process. Here, in addition to rewrite time, the monetary outlay for LLM inferencing also has to be considered.
7.3.1 Transformation Time
The average processing time per FPR query is shown in Table 9 for LITHE and SOTA over the various benchmarks. The good news is that the rewrites are typically identified in a few minutes even for highly complex queries. We hasten to note that this investment is likely to be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. As a case in point, for TPC-DS Q30 on a 10GB database, the original version took 17 minutes to run to completion, whereas the LITHE rewrite took 7 minutes to transform but finished in less than a second! Further, when the database size was increased to 100GB, the LITHE rewrite executed in just 8 seconds, while the original version took more than an hour. Finally, many applications tend to use a set of canned queries which are run thousands of times. Thus, a relatively small one-time investment can be easily recovered over repeat executions of such queries.
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down into the various components of LITHE is shown in Table 10 for the TPC-DS queries. We observe here that the lion’s-share is taken by the initial prompt ensemble and the final MCTS module, largely arising from their “brute-force” combinatorial construction. But both these modules can potentially be improved – the choice of prompt via a classifier, and MCTS by early pruning of ill-formed (semantic/syntactic errors) queries. These options are discussed below in Section 7.3.3.
Table 10: LITHE Transformation Time Analysis
Time/Query (min) Tokens / Query
Avg. Max. Avg. Max.
Basic Prompts (Section 3) 1.2 5.2 2407 17778
Rules R1 — R4 (Section 4.1) 0.8 3.6 3076 11988
Rules R5, R6 (Section 4.2) 0.5 2.1 2114 5865
MCTS (Section 5) 2.1 10.6 14792 55055
7.3.2 Monetary Outlay
The average number of LLM tokens required by LITHE and SOTA for the various benchmarks, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are shown in Table 9. The good news is that the inference charges per query are just a few cents, making rewriting practical from a deployment perspective.
We also see that LITHE can be a factor of two more expensive than SOTA. To understand the source of this expense, the token costs incurred by each module of LITHE are shown in Table 10. Again, the lion’s share is taken by the MCTS module. This is to be expected due to its exploratory nature which can result in traversal of several paths in the tree – however, the additional expense appears to be justified by its output of significantly better rewrites. Finally, even this expense gets reduced when the time overheads are cut down, as described below.
7.3.3 Overheads Reduction
Reducing the query rewrite time while still obtaining the same performance would be possible if we could directly use the MCTS-driven rewrite with the appropriate prompt. Towards this end, we build a classifier to pick the most appropriate rewrite rule for a given input query. Specifically, the classifier identifies which, if any, of Rules R1–R6 is appropriate for a given query. If none are appropriate, then it falls back to just the set of basic prompts to identify the best prompt to be given as input to the MCTS module. In addition to reducing the rewrite times, using a classifier can also reduce the financial costs of the rewrite.
We design an LLM based classifier to accomplish this as follows. The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied. For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision. Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 40 percent, and the tokens by about 20 percent. However, there is a price to be paid – the PR is reduced to 59 and the SpeedupGM comes down to 21.7. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
Pruning in MCTS. A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output. In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early. We are currently working on the design and implementation of such checks.
7.4 Alternative Platforms (DB/Engine/LLM/T2S)
7.4.1 Anonymized Database
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of the GPT-4o LLM having already been trained well on our evaluated benchmarks, all of which are in the public domain. To investigate this issue, we created an anonymized version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their content. We then constructed rewrites for the 65 FPR queries (after syntactic changes to reflect the new schema) on this version. The results are shown in Table 12 for SOTA and LITHE. We observe that while the SpeedupGM marginally decreases, the performance profiles remain similar to those seen earlier for native TPC-DS, testifying to the robustness of the rewrite approaches.
Table 11: Impact of Classifier (TPC-DS)
Metrics No Classifier With Classifier
# PR 65 59
SpeedupGM 30.6 21.7
Avg. Tokens 22390 18296
Avg. Time 4.6 min 2.8 min
Table 12: Rewrite Performance on Anonymized Database
Approach # PR SpeedupGM
TPC-DS Masked TPC-DS Masked
SOTA 16 14 2.1 1.8
LITHE 65 64 30.6 29.4
7.4.2 Commercial DBMS
Another legitimate question could be whether the rewrites made amends for the relative lack of sophistication in the PostgreSQL optimizer, and that its benefits would disappear in premier database engines. To evaluate this issue, we performed TPC-DS rewrites on two popular and highly engineered commercial DBMS.
Table 13 compares the performance of LITHE and SOTA on these systems. Compared to PostgreSQL, the PR reduced considerably from 65 to 15 and 29, respectively, in case of LITHE, while SOTA had just 6 and 14 PRs respectively. Nevertheless, a significant fraction of the benchmark queries are still represented, providing a very respectable SpeedupGM of 5.3 and 3.4. So, LITHE has a role to play even on high-end database engines. From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
Table 13: Rewrite performance on Commercial Database Engines
Approach # PR SpeedupGM
Opt A Opt B Opt A Opt B
LITHE 15 29 5.3 3.4
SOTA 6 14 2.2 2.5
7.4.3 LITHE on LLaMA
In our concluding experiment, we evaluate the performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which is said to have several hundred billion parameters [5]. To attain practical inference times, the model was loaded using a low 4-bit quantization. Further, to ensure reproducibility and deterministic answers, the do_sample parameter was set to False, which forces the LLM to perform greedy decoding. In order to make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
For this environment, Table 14 shows the PR and SpeedupGM obtained on the micro-benchmark with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (see Tables 2 and  5), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that even smaller models can be fruitfully used in real-world environments.
Table 14: Micro-benchmark Performance with LLaMA
# PR SpeedupGM
LLaMA without MCTS 5 2.6
LLaMA with MCTS 9 7.2
7.4.4 Text-to-SQL
As mentioned in the Introduction, we could alternatively construct the SQL-to-SQL rewrite problem as a two-stage pipeline: SQL-to-Text followed by Text-to-SQL. To assess this option, we first implemented SQL-to-Text on GPT-4o with the prompt asking for conversion of the SQL query to a clear business description. Then, two open tools were evaluated for Text-to-SQL – CHESS [28] and SQLCoder [3], which have performed very well on the BIRD [17] and SQL-Eval [2] benchmarks, respectively. However, in all our experiments, the outcomes were uniformly poor with the rewrites often suffering semantic errors, and even among the few correct rewrites, there was not a single PR. Therefore, this alternative architecture does not appear to be a viable option.
8 Related Work
Rule-based SQL rewriting. Most of the recent work on SQL query rewriting is rule-based [38, 31, 7, 32, 9, 22]. For instance, WeTune [31] employs a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules. While this approach can generate a large set of new rewrite rules, it often fails in coming up with transformation rules for complex queries due to the computational overheads of verifying rule correctness. As such, it is unable to rewrite any of the TPC-DS queries [14, 19].
Learned Rewrite [38] uses a set of existing rewrite rules from Calcite [8] and aims to learn the optimal subset of rules along with the order in which they must be applied. Given that the search space for possible rewrites is exponential in the number of applicable rules, an MCTS scheme is used to efficiently navigate this space and find the rewritten query with maximum cost reduction.
LLM-R2 [18] is also rule-based but takes a different approach to identify the order for rewrite rule applications: it uses an LLM to find the best Calcite rules and the order in which to apply them to improve the query performance. R-Bot [27] also leverages an LLM to optimize the order of Calcite rules, but employs advanced contemporary techniques such as retrieval-augmented generation (RAG) and step-by-step self-reflection to improve the outcomes.
Query Booster [7] implements human centred rewriting – it provides an interface that the user can use to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query. There also exist other rule-based rewrite approaches that are designed for specific type of rewrites such as optimizing correlated window aggregations [32] and common expression elimination [9].
All of the above approaches operate via the query plan space, which can restrict the kind of rewrites that can be accomplished. Whereas, LITHE uses a small set of general rewrite rules that work directly in the query space.
LLM-based rewriting. GenRewrite [19] is the first LLM-based approach that tries to use the LLM itself for end-to-end query rewriting. Instead of using predefined rules from Calcite [8], they employ the LLM to create Natural Language Rewrite Rules (NLR2s) to be used as hints, and perform several iterations of prompting to get the rewritten query. They show that LLMs can outperform rule-based approaches due to their ability to understand contexts, and demonstrate a significant improvement in query rewriting compared to prior methods.
A limitation, however, is that LLM-generated rewrite rules often fail to generalize beyond specific query pairs. Even when generalized rules are present, without accompanying examples, LLMs can struggle to apply rules correctly. Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
LLMs for Text-to-SQL. LLMs have also been extensively used for Text-to-SQL transformations [17, 33, 2, 3, 28]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [29, 23, 28]. On the other hand, the goal of SQL rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
9 Conclusions and Future Work
We investigated how the latent power of LLM technologies can be productively materialized in the context of SQL-to-SQL rewriting. Our study progressively infused database domain knowledge, such as redundancy removal rules and schematic+statistical metadata, into the LLM prompts. Further, the output telemetry of LLMs, in the form of token probabilities, was used to signal situations where the LLM lacked confidence, triggering exploration of a larger search space. Finally, a combination of logic-based and statistical tests was employed to verify the equivalence of the rewrites. An empirical evaluation over common database benchmarks showed that rewriting is a potent mechanism to improve query performance. The speedups obtained are significant and proof-of-concept, but there remains scope for improvement. For instance, while a speedup of 83X was produced for Q95 of TPC-DS, we were able to hand-craft an alternate rewrite with 99X speedup by moving a self-join from a CTE to the main query, thereby reducing scans of the web_sales table. We are currently working on adding a few select rules to implement such optimizations within the LITHE framework.
Our focus here was primarily on prompting-based strategies. In our future research, we also plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
Appendix
A1 : Detailed MCTS Algorithm
Algorithm 2 Token-augmented Rewrite
1:$root$,         # Start State;
2:$k$,            # Maximum number of child expansion
3:$\theta$,            # Probability threshold for node expansion
4:$iter_{max}$,     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
7:     $visits[root]$leftarrow $visits[root]$ + 1
8:     $n_{cur}$leftarrow $root$
9:     # Stage 1: Selection
10:     while $len(n_{cur}.children)>0$ do
11:         $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$
12:         $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
13:     # Stage 2: Expansion
14:     $expand\leftarrow$ True
15:     while $expand$ do
16:         $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
17:         if $tokens_{next}[0]=$‘;’ then
18:              $expand\leftarrow$ False
19:         else if $P_{next}\leq\theta$ then
20:              for $token\in tokens_{next}$ do
21:                  $state\leftarrow n_{cur}.state\cdot token$
22:                  $n_{new}\leftarrow$ new $Node$ with State $state$
23:                  Append $n_{new}$ to $n_{cur}.children$
24:              $expand\leftarrow$ False
25:         else
26:              $state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$ First Token
27:              $n_{new}\leftarrow$ new $Node$ with State $state$
28:              Append $n_{new}$ to $n_{cur}.children$
29:              $n_{cur}\leftarrow n_{new}$
30:     # Stage 3: Simulation
31:     $query\leftarrow$ GreedyExpand($n_{cur}$)
32:     # Compute the potential of query (returns 0 for invalid query)
33:     $v\leftarrow$ ComputePotential($query$)
34:     $Potential[query]\leftarrow v$
35:     # Stage 4: Backpropagation
36:     $n\leftarrow n_{cur}$
37:     while $n\neq$ Null do
38:         if $v>V[n]$ then
39:              $V[n]\leftarrow v$
40:         $n\leftarrow\text{Parent}(n)$
41:# Return valid rewrite with maximum $Potential>1$
42:if $\exists q\in Potential\mid Potential[q]>1$ then
43:     return $q$ having maximum value of $Potential[q]$
44:else
45:     return the original query
A2 : Query Equivalence Testing
We use a multi-stage approach, described below, to test semantic equivalence between the original query and a candidate rewrite.
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g. Cosette[10], SQL-Solver [13], QED [30]). In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
To probabilistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples. The goal is to reduce the likelihood of non-equivalent queries returning the same results.
To minimize the occurrence of the second problem (empty results), the following approach is taken:
  1. 1.
    We use correlated sampling [34] to sample the database. This technique leverages the join graph of the schema to produce a sample that maintains join integrity between the tables participating in the query.
  2. 2.
    To differentiate between inner and outer joins, we insert into the sample database, rows with appropriate values on the relevant columns such that they produce NULL values for outer joins.
  3. 3.
    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result. We make use of the rows in the sampled data for this purpose. For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database. We then replace the query constant with a value already present in the sample. Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database. Note that these modifications only change the selectivity of the query, but not its semantics.
3. Result Equivalence on the Entire Database. Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself. However, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
A3 : Examples used in Rules 1–6
R1: Use CTEs (Common Table Expressions) to avoid repeated computation.
Original Query
SELECT emp.employee_name,
       mgr.manager_name
FROM   employees emp,
       managers mgr
WHERE  emp.manager_id = mgr.manager_id
       AND emp.employee_id IN (SELECT manager_id
                               FROM   (SELECT manager_id,
                                              manager_name
                                       FROM   managers
                                       WHERE  job_id = ’IT_PROG’
                                              AND manager_id > 100))
       AND mgr.manager_name IN (SELECT manager_name
                                FROM   (SELECT manager_id,
                                               manager_name
                                        FROM   managers
                                        WHERE  job_id = ’IT_PROG’
                                               AND manager_id > 100));
Rewritten Query
WITH cte
     AS (SELECT manager_id,
                manager_name
         FROM   managers
         WHERE  job_id = ’IT_PROG’
                AND manager_id > 100)
SELECT emp.employee_name,
       mgr.manager_name
FROM   employees emp,
       managers mgr
WHERE  emp.manager_id = mgr.manager_id
       AND emp.employee_id IN (SELECT manager_id
                               FROM   it_prog_managers)
       AND mgr.manager_name IN (SELECT manager_name
                                FROM   it_prog_managers);
R2: When multiple subqueries use the same base table, rewrite to scan the base table only once.
Original Query
SELECT (SELECT Avg(salary)
        FROM   employees
        WHERE  department = ’Sales’
               AND experience_years BETWEEN 1 AND 5
               AND salary BETWEEN 50000 AND 60000) AS Sales_Avg,
       (SELECT Avg(salary)
        FROM   employees
        WHERE  department = ’HR’
               AND experience_years BETWEEN 5 AND 10
               AND salary BETWEEN 80000 AND 90000) AS HR_Avg;
Rewritten Query
SELECT avg(
       CASE
              WHEN department = ’Sales’ THEN salary) AS sales_avg,
       avg(
       CASE
              WHEN department = ’HR’ THEN salary) AS hr_avg
FROM   employees
WHERE  (
              department = ’Sales’
       AND    experience_years BETWEEN 1 AND    5
       AND    salary BETWEEN 50000 AND    60000)
OR     (
              department = ’HR’
       AND    experience_years BETWEEN 5 AND    10
       AND    salary BETWEEN 80000 AND    90000);
R3: Eliminate overlapping subqueries.
Original Query
SELECT c.*
FROM   customer c
WHERE  c.address_id IN (SELECT a.address_id
                        FROM   address)
       AND c.address_id IN (SELECT a.address_id
                            FROM   address
                            WHERE  a.pin_code = ’560012’);
Rewritten Query
SELECT c.*
FROM   customer c
WHERE  c.address_id IN (SELECT a.address_id
                        FROM   address
                        WHERE  a.pin_code = ’560012’);
R4: Remove unnecessary joins between a primary key and a foreign key.
Schema
CREATE TABLE products
     p_product_id INTEGER NOT NULL,
     PRIMARY KEY (p_product_id)
CREATE TABLE fact_sales
     f_sales_id   INTEGER NOT NULL,
     f_units_sold INTEGER NOT NULL,
     f_product_id INTEGER NOT NULL,
     PRIMARY KEY (f_sales_id),
     FOREIGN KEY (f_product_id) REFERENCES products(p_product_id)
Original Query
SELECT p_product_id,
       f_units_sold
FROM   fact_sales,
       products
WHERE  f_product_id = p_product_id;
Rewritten Query
SELECT f_product_id,
       f_units_sold
FROM   fact_sales;
R5: Choose EXIST or IN based on subquery selectivity.
Original Query
Select item.id, item.code, item.price
from item
where item.sourceid in (
    Select element.sourceid
    from element
    where element.zip > 1100
order by item.id;
Statistics
Selectivity of different predicates is given below :
( 1 ) source_id > 1100 on table element :: 0.7385
Rewritten Query
Select item.id, item.code, item.price
from item
where exists(select 1
    from element
    where item.sourceid = element.sourceid
      and element.sourceid > 1100
order by item.id;
R6: Pre-filter fact tables in a CTE using those dimension tables having low selectivity values. Retain dimension table filters in main query. Do not create explicit join statements.
Original Query
Select item.id, item.code, item.price, country.name
from item, element, country, colour
where element.sourceid = item.sourceid
  and item.code = country.code
  and item.colorcode = colour.colorcode
  and element.zip > 8000
  and country.name = ’US’
  and colour.colour_name = ’Red’
order by item.id;
Statistics
Selectivity of different predicates is given below :
( 1 ) element.zip > 8000 :: 0.002
( 2 ) country.name = ’US’ :: 0.4
( 3 ) colour.colour_name = ’Red’ :: 0.01
Rewritten Query
with filtered_item as(
  Select *
  from item
  where
    item.sourceid in (
      Select
        element.sourceid
      where
        element.zip > 8000
    and item.colorcode in (
      Select
        colour.colorcode
      from
        colour
      where
        colour.colour_name = ’Red’
Select item.id, item.code, item.price, country.name
from filtered_item as item, element, country, colour
where
  item.sourceid = element.sourceid
  and item.code = country.code
  and item.colorcode = colour.colorcode
  and element.zip > 8000
  and country.name = ’US’
  and colour.colour_name = ’Red’
order by item.id;

### v2

@@
 Abstract
-Query rewriting is a classical technique for transforming complex declarative SQL queries into “lean” equivalents that are conducive to (a) faster execution from a performance perspective, and (b) better understanding from a developer perspective.
-The rewriting is typically achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system.
-In recent times, LLM-based techniques have also been mooted, but they are prone to both semantic and syntactic errors.
-We investigate here, how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency.
+When complex SQL queries suffer slow executions despite query optimization, DBAs typically invoke automated query rewriting tools to recommend “lean” equivalents that are conducive to faster execution.
+The rewritings are usually achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system.
+Recently, LLM-based techniques have also been suggested, but they are prone to semantic and syntactic errors.
+We investigate here how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency.
 Our study shows that these goals can be progressively achieved through incorporation of (a) an ensemble suite of basic prompts, (b) database-sensitive prompts via redundancy removal and selectivity-based rewriting rules, and (c) LLM token probability-guided rewrite paths.
-Further, a suite of statistical and logic-based tools can be used to guard against errors produced by the model.
-We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from multiple benchmarks on contemporary database platforms.
-The results show significant improvements over SOTA rewriting techniques – for instance, on TPC-DS, LITHE constructed productive ($>$1.5x speedup) rewrites for two-thirds of the query suite, delivering four times more coverage than SOTA.
-Further, the geometric mean of its estimated execution speedups was an order-of-magnitude jump over SOTA performance.
-In essence, LITHE offers a potent and robust LLM-based intermediary between enterprise applications and database engines.
+Further, a suite of logic-based and statistical tools can be used to check for semantic violations in the rewrites prior to DBA consideration.
+We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms.
+The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer.
+For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
+Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise query performance.
 1 Introduction
 The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools.
-A compelling case in point is the blog-processing query [4] shown in Figure 2 – this query was produced by the popular Entity Framework [21] mapper.
-It computes a daily summary of rating metrics for highly-rated blogs.
-This seemingly complex nested query can be equivalently rewritten in the “lean” flat version shown in Figure 2.
+Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans.
+However, in practice, they are often led astray by complex representations, resulting in poor response times.
 SELECT t.Key, sum(t.Rating) AS PostRating,
@@
 b1.Url, p0.day AS Key
-FROM Posts AS p0 INNER JOIN Blogs AS b1 ON
-p0.BlogId = b1.BlogId
+FROM Posts AS p0 INNER JOIN Blogs AS b1
+ON p0.BlogId = b1.BlogId
 WHERE b1.Rating > 5) AS t0
@@
 Figure 2: Lean Equivalent Query
-Albeit less common, we observe such superfluity even in the hand-crafted expert query domain – for instance, TPC-DS Query Q23 includes a sub-query that identifies customers with the maximum store sales in a four-year period; this sub-query has a redundant join of store_sales with customer despite the FK-PK relationship between these tables, as highlighted in [25].
-Transforming bloated queries to lean equivalents in the application source code is beneficial from multiple perspectives: First, rewrites can improve code readability, a major concern in industrial contexts.
-Second, query optimizers are expected to, in principle, remove bloat while constructing efficient execution plans.
-However, in practice, they are often led astray by complex representations, resulting in poor query performance.
-In fact, both examples discussed above were implemented as such by the current PostgreSQL optimizer (v16) [1], without any change.
-Therefore, a query rewriter can serve as an effective and non-invasive mechanism for delivering good performance despite inherent optimizer limitations.
-Given this motivational context, a viable SQL-to-SQL query transformer should satisfy the following criteria: (1) The transformed query should be semantically equivalent to the original; (2) There should ideally be performance improvement due to the rewrite, but at the very least, no regression in query performance; and (3) The transformation overheads must be practical for deployment.
+As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs.
+When this seemingly complex query is given to the current PostgreSQL engine (v16) [1], its execution plan essentially mimics the hierarchical structure of the query.
+This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete.
+However, the query can be equivalently rewritten (assuming NOT NULL column constraints) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
+For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives.
+For this context, a viable S2S query transformer should satisfy the following criteria: (1) A wide range of queries should be covered in its scope; (2) The transformed query should be semantically equivalent to the original; (3) The rewrite should ideally improve performance, but at least not cause regression; and (4) The transformation overheads must be practical for deployment.
 Prior Work
@@
 Learned Rewrite [45]) and Model-based (e.g.
-Gen-Rewrite [19]) approaches have been proposed in the literature for SQL query transformations.
-However, as explained later in the related work section (Section 8), the benefits of these state-of-the-art (SOTA) techniques are curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space.
-An alternative and extremely simple strategy would be to take advantage of the large body of recent literature on LLM-based Text-to-SQL translations (e.g.
-[17, 35, 33]).
-That is, first invoke an LLM to transform the original SQL to text, and then employ a Text-to-SQL tool to produce quality SQL.
-However, as discussed later in Section 7, we found this approach to be highly prone to semantic errors and lacking execution benefits.
-This is perhaps to be expected given the focus of Text-to-SQL tools on query formulation, orthogonal to our objective of performance enhancement.
-Therefore, the SQL-to-SQL problem merits a fresh assessment from first principles.
+Gen-Rewrite [25]) approaches have been proposed in the literature for S2S transformations.
+Further, a recent vision paper [46] has persuasively advocated the use of LLMs for such query rewriting.
+These novel state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting.
+However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space – the latter provides greater scope for candidate rewrites since only intent, and not implementation, is expressed.
+In this paper, we address these lacuna and amplify the effectiveness of query rewriting.
+Figure 3: High-level architecture of LITHE
 The LITHE Rewriter
-In this study, we carry out a calibrated investigation of how LLMs could be reliably used to produce performant rewrites of complex SQL queries.
-The proposed techniques, described below, have been implemented and evaluated in the LITHE (LLM Infused Transformations of Hefty queries) rewriting system developed by our group.
+We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention.
+As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance.
+Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$.
+Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
 Basic Prompts.
-We begin with an ensemble suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs.
-Interestingly, we find that more information is not necessarily better wrt rewriting quality, and the best prompt granularity is query-specific.
-Moreover, this simple prompt ensemble was found to deliver performance similar to the SOTA techniques.
+We begin with a suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs.
+Interestingly, we find that more information is not necessarily better wrt rewriting quality, and that the best prompt granularity is query-specific.
+Moreover, this basic prompt ensemble was found to itself deliver performance similar to the SOTA techniques.
 Database-sensitive Prompts.
 To help the model adapt to different query patterns and structures, we next introduce rules in the prompts.
-But our choice and usage of rules are markedly different from those used in prior work.
-First, our rules are invoked directly in query space.
-The LLM therefore gains the latitude to generalize the rule usage to a wide range of queries.
+Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries.
 This is in contrast to the hardwired and narrow rule application mechanisms (e.g.
 Calcite [9] rules) typically used in existing rewrite systems, which operate in plan space.
-Second, to minimize rule application overheads, we incorporate only a handful of rules that serve to eliminate the primary causes of poor query performance.
-Whereas current rewrite systems employ a large number of rules in their processing.
 In particular, we work with two classes of rules:
@@
 Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts.
-To our knowledge, such metadata inclusion has not been considered before in the SQL rewrite context.
+To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context.
 As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
@@
 MCTS has previously demonstrated significant effectiveness in code generation exercises [43].
+Semantic Equivalence Labels.
+After generating a rewrite, it is imperative to check for semantic equivalence to the original query.
+Unfortunately, the general problem of proving query equivalence is NP-complete [6].
+However, for restricted classes of queries, various logic-based equivalence-proving tools are available, and in LITHE, we use the recently developed QED tool [36] to verify rewrite correctness for queries within its scope.
+The remaining out-of-scope queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database.
+Finally, result equivalence can be optionally tested on the original database itself.
+While these checks do not suffer from false negatives, they may (rarely) incur false positives.
+After carrying out the above exercise, LITHE provides a verification label – provable or statistical – regarding semantic equivalence.
+Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
 Experimental Evaluation
-Our experiments are carried out on a variety of synthetic and real benchmarks: TPC-DS [11], DSB [12], JOB [15], StackOverflow [20] and Archer [37], covering a broad spectrum of databases and queries.
-Further, anonymized versions, where the table and column names are altered to not convey any semantic meaning about their contents, are also evaluated to assess robustness to databases not seen by the LLM during its training phase.
-These databases are hosted on PostgreSQL and GPT-4o is used as the LLM.
-We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [38], LLM-R2 [18], GenRewrite [19]) as well as a baseline LLM prompt [19].
-The primary metrics are the execution cost reductions and the rewriting overheads.
-For LLM-based techniques, the number of tokens in the prompts is also monitored since the financial charges for LLM usage are typically dependent on this number.
-To understand the impact of the various components of LITHE, a systematic ablation study is carried out.
-Finally, the performance on alternative database and LLM platforms is also evaluated.
-Verifying Semantic Equivalence.
-The general problem of proving query equivalence has long been known to be computationally hard [6].
-However, for restricted classes of queries, various logic-based equivalence-proving tools are available (e.g.
-Cosette [10], SQL-Solver [13], and QED [30]).
-In our LITHE system, we use QED [30] to verify rewrite correctness for queries within its scope.
-The remainder are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases using small sampled versions of the original database.
-We also provide the user with an option to test result equivalence using the original database.
+Our first set of experiments to evaluate LITHE’s performance is carried out on the industry standard TPC-DS benchmark [12], hosted on the PostgreSQL platform with GPT-4o used as the LLM.
+The evaluation focuses on slow queries taking more than a threshold time to complete.
+We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [45], LLM-R2 [24], GenRewrite [25], as well as a baseline LLM prompt [25]).
+The primary metrics are (a) reductions in optimizer-estimated costs, (b) run-time speedups, and (c) rewriting overheads.
+For LLM-based techniques, the number of tokens used is also monitored since the financial charges for LLM usage are typically dependent on this number.
+To understand the independent utility of the various components of LITHE, a systematic ablation study is carried out.
+Generalizability.
+In our second stage of experiments, we evaluate generalizability of the above outcomes in a variety of new scenarios, including (a) commercial database engines, (b) unseen database schemas, and (c) alternative LLM platforms.
 Results.
-Our experiments demonstrate that LITHE achieves semantically correct transformations that deliver a significant reduction in estimated execution times for many queries.
-Moreover, LITHE performed better than or at least as well as SOTA on all the benchmarks.
-As a compelling case in point, LITHE constructed “productive” (¿1.5x speedup) rewrites for about two-thirds of the complete TPC-DS query suite, delivering four times more coverage than SOTA techniques.
-Further, the GM (Geometric Mean) of the reduction in query execution costs was 30.6, an order-of-magnitude jump over that offered by SOTA.
-In essence, LITHE offers a potent LLM-based intermediary between enterprise applications and database engines.
+Our experiments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract costs.
+In particular, for TPC-DS, LITHE constructed “highly productive” ($>1.5x$ estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number.
+Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
+We also evaluated whether the above cost reductions translated to real execution speedups.
+Here, we find that LITHE is indeed often substantively faster at run-time as well.
+Specifically, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
+However, we also found that the well-known discrepancies between optimizer cost predictions and actual execution times [26] could be exacerbated by an injudicious choice of rule formulations.
+This separation suggests a future line of research to reduce the reasoning distance between language-based foundation models and statistics-based optimizer models.
+Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
 Contributions
@@
   1. 1.
-    Assesses LLM suitability for SQL-to-SQL transformation.
+    Assesses LLM suitability for S2S transformation.
   2. 2.
-    Transforms directly in query space instead of a plan space intermediate representation, leading to robust rewrites.
+    Transforms directly in query space instead of plan space intermediates, leading to performant rewrites.
   3. 3.
-    Incorporates potent database-sensitive rules in LLM prompts covering schematic and statistical dimensions.
+    Incorporates database-sensitive rules in LLM prompts, covering both schematic and statistical dimensions.
   4. 4.
@@
   5. 5.
-    Evaluates rewriting robustness over a broad range of database environments and benchmarks.
+    Evaluates rewrite quality over a broad range of database environments, demonstrating substantial benefits over both SOTA and the native optimizer.
   6. 6.
-    Demonstrates substantive performance benefits, well beyond SOTA, from LLM-powered query rewrites.
+    Identifies learnings that could help guide research directions for industrial-strength query rewriting.
 2 LITHE Overview
-The LITHE architecture, illustrated in Figure 3, consists of two main components: A prompt-based rewriting pipeline, and a token probability-driven rewrite exploration.
-Prompt-based rewriting pipeline.
-This pipeline comprises four modules: (1) LLM Prompting, (2) Syntax Verification, (3) Query Costing, and (4) Semantic Verification.
+The LITHE architecture, illustrated in Figure 3, consists of the following five-stage pipeline:
+1.
+Prompt-based rewriting.
+This component consists of two modules: LLM Prompting and Syntax Verification.
 The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite.
-The LLM output is checked by the syntax verification component – if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively until a valid SQL query is obtained.
-Then, the execution cost of this version is evaluated – if more than the cost of the original query, the rewrite is discarded.
-Otherwise, the rewrite is tested for semantic equivalence.
-If a match is established, the rewrite is returned; otherwise, the original query itself is returned.
-Finally, to prevent runaway situations, a threshold is set on the number of syntax correction attempts (5 in our experiments) and the original query is returned if this threshold is breached.
-The above process is repeated for each of the different prompts described in Sections 3 and 4.
-Then, the prompt providing the least-cost rewrite is returned.
-Since each prompt is evaluated within a fresh LLM context, there is no dependency on the order in which they are assessed.
-Within the pipeline, LITHE uses (a) the query parser from the database engine for syntax verification, (b) the query optimizer for cost estimation, and (c) the techniques described in Section 6.3 to check semantic equivalence.
-Figure 3: High-level architecture of LITHE
+The LLM output’s syntax is then verified with the database engine’s parser– if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively (within a threshold) until a valid SQL query is obtained.
+This process is repeated for each of the different prompts described in Sections 3 and 4, and the resulting syntactically-valid rewrites are fed to the query costing component.
+We evaluate each prompt within a fresh LLM context, thereby making the prompt processing to be order-independent.
+2.
+Query Costing.
+The costs of candidate rewrites are evaluated via the database engine’s optimizer.
+Rewrites whose costs are greater than that of the original query are immediately discarded.
+Whereas, the potentially beneficial rewrites (if any) are checked for semantic equivalence to the original query.
+3.
+Semantic Equivalence.
+A suite of provable (logic-based) and statistical (result-based) techniques are employed to assess the semantic equivalence of a recommended rewrite.
+If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
+4.
 Token probability-driven rewrite.
-The pipeline-chosen prompt is used as an input to a Monte-Carlo tree search (MCTS)-based procedure to further improve the query rewrite quality.
+The prompt producing the most beneficial (and valid) rewrite is used as input to a Monte-Carlo tree search (MCTS)-based procedure to further refine the rewrite quality.
 Specifically, multiple exploration paths are followed whenever the LLM lacks high confidence in the predicted token (details in Section 5).
+The query costing and semantic equivalence modules are also used internally within this procedure.
+5.
+Output.
+The least expensive valid rewrite as identified by the MCTS module is returned to the DBA together with the cost of the rewritten query and the approach used to claim equivalence.
+Additionally, an LLM-generated reasoning for the expected performance improvement is also provided.
+If no valid rewrite is identified, the original query itself is returned.
 2.1 Performance Framework
-As mentioned in the Introduction, the majority of our experiments are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine.
-We define a productive rewrite (PR) as one that improves the query’s performance by at least 1.5 times wrt the execution times estimated by the database engine’s query optimizer – this aggressive choice of threshold is to make the rewrite outcome desirable from an industrial perspective.
-Given a query workload, the set of feasible productive rewrites (FPR) consists of the queries for which PR are known to be feasible by some technique or the other.
-Note that it may be inherently impossible to produce PR on queries that are structurally simple or efficiently written by human experts.
-Finally, a maximum productive rewrite (MPR) is when the PR speedup of the rewrite matches the best known PR on this query.
-The overall benefit provided by a rewriting tool is quantified by the number of PR obtained on the query workload.
-Additionally, we also measure SpeedupGM, the geometric mean (GM) of the speedups obtained on the FPR queries in this workload.
-(The geometric mean is used to ensure equal weight for all queries in the workload, irrespective of their relative durations.)
+We consider a query that takes more than T seconds to complete on the native database engine as a “slow query”, potentially triggering intervention by the DBA.
+Based on common industry perception (e.g.
+[4]), T is set at 10 seconds in our study.
+For this context, we define a Cost Productive Rewrite (CPR) as a rewrite that improves a slow query’s performance by at least 1.5 times wrt the optimizer-estimated cost – this aggressive choice of threshold is so that: (a) there is enough headroom in the optimizer prediction that a runtime regression is unlikely; and (b) the benefits of the rewrite substantively outweigh the transformation overheads.
+The overall benefit provided by a rewriting tool is quantified by the number of CPR obtained on the slow query workload.
+Additionally, we also measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites.
+Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites.
 2.2 Query Micro-benchmark
-To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft productive rewrites.
-These human rewrites deliver a SpeedupGM of 13.9 on the micro-benchmark, serving as an informal “upper-bound” on what is computationally attainable.
-Later, in Section 7, we extend the evaluation to a suite of complete benchmarks.
+To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft high-quality CPRs.
+These queries are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine.
+The human rewrites deliver a CSGM of 11.84, serving as an aspirational target to attain computationally.
+Later, in Section 7, we extend the evaluation to complete benchmarks.
+Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sec 345). The TSGM performance is subsequently discussed in Section 7.
 3 Basic Prompts
 In this section, we explore the simplest interface to LLMs, namely prompting, for query rewriting.
-We evaluate four basic prompts, illustrated in Figure 4, which cover a progressive range of detail in the instructions and test the effectiveness of the LLM’s base knowledge.
+We evaluate four basic prompts, enumerated in Figure 4, which cover a progressive range of instructional detail and test the effectiveness of the LLM’s base knowledge.
 Figure 4: Templates used for Basic Prompts
@@
 Finally, it is instructed to apply the identified solution.
-Essentially, the prompt tries to make the LLM reason akin to a human expert.
+Essentially, the prompt tries to make the LLM reason akin to human experts.
 Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once.
@@
 Performance
-The performance of the four prompt templates on the micro-benchmark (where all 10 queries are in FPR) is shown in Table 1.
+The performance of the four prompt templates on the micro-benchmark is shown in Table 1.
 We find that less than half the rewrites are productive with individual prompts.
-However, a drill-down shows that each prompt does the best on certain queries – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them.
-While this ensemble approach (Row 5 in Table 1) raises the PR to 6, there remain four queries that could not be productively rewritten by the prompts alone.
-Table 1: Performance of Basic Prompts
-Prompt # PR SpeedupGM
-Prompt 1 3 2.7
-Prompt 2 4 1.57
-Prompt 3 3 1.37
-Prompt 4 2 1.27
-Prompt Ensemble 6 3.6
-SOTA Ensemble 5 3.4
-The SpeedupGM, shown in the last column, for each individual prompt is less than 3, while that of the ensemble reaches 3.6.
-But these speedups, while productive, are all lower than those of the human rewrites, that is, they are not MPR.
-Finally, as another baseline, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform.
-They delivered (Row 6 in Table 1) 5 productive rewrites with a GM of 3.4, indicating the wide gap between the current state-of-the-art and what is possible.
+However, a drill-down shows that the best prompt in the ensemble is query-specific – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them.
+This ensemble approach raises the CPRs to 6 (Row 5 in Table 1) – however, there remain four queries that are not productively rewritten by these prompts.
+Table 1: Performance of Basic Prompts.
+Prompt # CPR CSGM
+Prompt 1 3 1.99
+Prompt 2 2 1.85
+Prompt 3 4 2.83
+Prompt 4 4 3.00
+Prompt Ensemble 6 3.23
+SOTA Ensemble 3 2.49
+The CSGM, shown in the last column of Table 1, is at most 3 for the individual prompts, while the ensemble reaches 3.23.
+But these speedups, although productive, are all lower than those delivered by the human rewrites.
+Finally, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform.
+They delivered 3 CPRs with a CSGM of 2.49 (last row in Table 1), indicating the wide gap between the current reality and what is humanly possible.
 4 Database-Sensitive Prompts
@@
 Rules R1 through R4 in Table 2 are designed to tackle such redundancies.
-For instance, the redundant PK-FK join highlighted in the Introduction is addressed by rule R4 in Table 2 (the relevant schematic information is also provided in the prompt).
-Table 2: Rules for Database-sensitive prompts
+The relevant schematic information (e.g.
+table names, column names, constraints) required by these rules is also provided in the prompt.
+Table 2: Rules for Database-sensitive prompts.
 Redundancy Removal Rules
-R1 Use CTEs (Common Table Expressions) to avoid repeated computation.
+R1 Use CTEs (Common Table Expressions) to avoid repeated computation of a given expression.
 R2 When multiple subqueries use the same base table, rewrite to scan the base table only once.
@@
 R5 Choose EXIST or IN from subquery selectivity (high/low).
-R6 Pre-filter fact tables in a CTE using dimension tables with low selectivities. Retain dimension table filters in main query. Do not create explicit join statements.
-The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are enumerated in the Appendix A3.
+R6 Pre-filter tables involved in self-joins and with low selectivities on their filter and/or join predicates. Remove any redundant filters from the main query. Do not create explicit join statements.
+Figure 5: Templates for Database-sensitive Prompts
+The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are available in the Section E.
 Note that this prompt template allows for only a single rule to be present in the prompt.
@@
 Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
-Figure 5: Templates for Database-sensitive Prompts
 Performance
 The performance improvement achieved on the micro-benchmark by an ensemble that adds the redundancy-removing prompts to the basic set (Section 3) is shown in Table 3.
-We observe that the PR increases to 8, and SpeedupGM grows to 4.8.
-Table 3: Performance with Redundancy Removal Rules
-Prompt # PR SpeedupGM
+We observe that the CPR increases to 7, and CSGM grows to 6.85.
+Table 3: Performance with Redundancy Removal Rules.
+Prompt # CPR CSGM
 […]
@@
 For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values.
-Based on such arguments, rules R5 and R6, which pre-filters fact tables for dimension tables with low selectivity values, are included in Table 2.
-Note that, a specific instruction to not create explicit joins had to be added in rule R6.
+Rule R5 basically encodes this argument as a rule in Table 2.
+Similarly rule R6, which pre-filters tables that are involved in self-joins and have low selectivity filter and/or join predicates, is also used to guide a rewrite.
+Note that a specific instruction to not create explicit joins had to be added in rule R6.
 This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@
 The performance improvements following addition of selectivity-guided prompts are shown in Table 4.
-We observe that PR are now obtained for all 10 micro-benchmark queries.
-Moreover, the resulting SpeedupGM increases to 11.8, quite close to the human upper bound of 13.9.
-Table 4: Performance of Metadata-infused Prompts
-Prompt # PR SpeedupGM
+We observe that CPRs are now obtained for 9 of the ten micro-benchmark queries.
+Moreover, the resulting CSGM increases to 10.57, quite close to the human target of 11.84.
+Table 4: Performance of Metadata-infused Prompts.
+Prompt # CPR CSGM
 […]
-A key point to note in closing is that rules R1 through R6 not only bring additional queries into the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3.
-Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule, and this option is examined in Section 7.3.3.
+We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3.
+Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in the Section B.
 5 Token Probability Driven Rewrite
 A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up.
-This is often due to the potential output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs.
+This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs.
 In order to have a robust approach for such cases, we take inspiration from the code generation literature [43].
@@
 The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query.
-Given that the vocabulary sizes of LLMs are upward of hundreds of thousands of tokens, it may become very expensive (in both financial and computational costs) to construct the entire tree.
+Given the hundreds of thousands of tokens in LLM vocabularies, it may be very expensive (both financially and computationally) to construct the entire tree.
 It is therefore essential to significantly reduce the token search space while exploring the tree for valid rewrites.
 This is precisely the purpose of MCTS which applies an Upper Confidence Bound (UCB) heuristic [31] to identify the best paths in a tree without computing the entire tree.
-5.1 MCTS Search Process
-The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
-1.
-Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure).
-Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1).
-Actions are picked using a UCB that balances exploration and exploitation.
-The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
-Specifically, given a node $n$ and a set of possible actions $a\in A$, the next node in this traversal is chosen as:
-$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
-where $UCB$ is a heuristic sourced from [26], modified to reflect our formulation where values are associated with nodes rather than edges in the tree. It is defined as follows:
-$UCB(n,a)=V(n^{\prime})+\beta(n)*P_{LLM}(a|n.state)*\frac{\sqrt{log(visits[n])}}{1+visits[n^{\prime}]}$ (2)
-Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3).
-The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often.
-Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process.
-$\beta$ is a function that controls the balance between exploration and exploitation.
-It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore.
-$\beta$ is defined as:
-$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
 Algorithm 1 Token-augmented Rewrite
@@
 36:    return the original query
+5.1 MCTS Search Process
+The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
+1.
+Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure).
+Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1).
+Actions are picked using a UCB that balances exploration and exploitation.
+The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
+Specifically, given a node $n$ and a set of possible actions $a\in A$, the next node in this traversal is chosen as:
+$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
+where $UCB$ is a heuristic sourced from [31], modified to reflect our formulation where values are associated with nodes rather than edges in the tree. It is defined as follows:
+$UCB(n,a)=V(n^{\prime})+\beta(n)*P_{LLM}(a|n.state)*\frac{\sqrt{log(visits[n])}}{1+visits[n^{\prime}]}$ (2)
+Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3).
+The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often.
+Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process.
+$\beta$ is a function that controls the balance between exploration and exploitation.
+It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore.
+$\beta$ is defined as:
+$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
 2.
@@
 This, however, is expensive both from the aspect of query rewrite time, as well as the number of LLM tokens used.
-To minimize these costs, the LITHE workflow first selects, given a query, the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4.
+To minimize these costs, given a user query, the LITHE workflow first selects the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4.
 It then employs this prompt to initiate the MCTS-based rewrite.
 In case no prompt provides a lower-cost rewrite, baseline Prompt 1 of Section 3 is used as the fallback option.
-Later, in Section 7.3.3, we show how the overheads incurred can be further reduced via a classifier construction.
 5.3 Performance
-The performance results with MCTS-based rewrites are shown in Table 5 for the micro-benchmark. We observe gratifyingly that the SpeedupGM now attains the human speedup of 13.9.
-Table 5: Performance of MCTS-based rewrites
-# PR SpeedupGM
-LITHE 10 13.9
+We find that LITHE’s performance with MCTS-based rewrites fully matches the human target, providing 10 CPRs and the maximum CSGM of 11.84.
+While these improvements, as compared to just prompting, might seem marginal at this stage, we wish to make two observations: (1) MCTS helps extract the maximum improvement wherever possible; and (2) As shown later in Section 7.5.2, improvements due to MCTS become significant for smaller LLMs such as LLaMA.
 6 Implementation Choices
@@
 6.1 LITHE Parameter Settings
-The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response, with a higher temperature resulting in the outputs being more varied and less predictable.
+The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response.
 While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible.
-Hence we set the temperature to 0 which causes the model to do greedy sampling to select the next token.
+Hence we set the temperature to 0 which forces the model to greedily sample the next token.
 The hyperparameters used by LITHE for MCTS are as follows: The maximum number of iterations $iter_{max}$ is set to 8, expansion threshold $\theta$ is 0.7, and number of expansions $k$ is 2.
@@
 These settings were determined after an empirical evaluation of the various trade-offs, providing a robust balance between efficiency and quality.
-Finally, we run the prompt pipeline for a maximum of 5 iterations to try and fix any syntax errors created during the rewrite process.
-6.2 Query Efficiency Characterization
-The LITHE workflow requires computing the performance of several candidate query rewrites.
-In this scenario, explicitly executing each rewrite on the database to evaluate its runtime benefit may become impractically slow.
-This is especially true for queries where even a single execution is time-consuming.
-We therefore leverage the query cost, as estimated by the optimizer, to evaluate the performance of query rewrites.
-It is common knowledge that these estimates are often inaccurate [16] – however, they are expected to be generally reflective of the actual run times, especially on modern industrial-strength database systems.
-Moreover, our comparison of costs is only in a relative sense – between the original query and its rewrite, so errors in the absolute values are tolerable as long as similar errors are made in all estimates.
-These costs can therefore serve as a computationally cheap and meaningful measure to distinguish performance among rewrites.
+Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2).
 6.2 Query Equivalence Testing
@@
 Logic-based Equivalence.
-Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g.
-Cosette[10], SQL-Solver [13], QED [30]).
-In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives.
+Although verifying the equivalence of a general pair of SQL queries is NP-complete [6], a variety of logic-based tools (e.g.
+Cosette[11], SQL-Solver [15], VeriEQL [17], QED [36]) are available for proving equivalence over restricted classes of queries, as mentioned in the Introduction.
+In LITHE, we use the recently proposed QED [36] since it was found to cover a larger set of queries compared to the alternatives.
 The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
@@
 If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence.
-The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results.
+The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results.
 However, while this test is a necessary condition for query equivalence, it is not sufficient.
 That is, false positives may be present because the sampled database may not cover all the predicates featured in the query.
-To minimize this likelihood, we use a combination of (1) correlated sampling [34] for maintaining join integrity in the sample, (2) addition of synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Appendix A2.
+To minimize this likelihood, we use a combination of (1) correlated sampling [42] for maintaining join integrity in the sample, (2) adding synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Section C.
 3.
 Result Equivalence on the Entire Database.
-Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself.
-However, as mentioned above, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions.
-Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
+Result equivalence could also be evaluated, in principle, on the entire database itself.
+Of course, this could turn out to be prohibitively expensive, especially if the queries themselves are time-consuming (e.g.
+due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions.
+Therefore, we leave this check as an optional choice for the DBA.
 7 Experimental Evaluation
@@
 Then we present the speedup results for both aggregate benchmark and individual queries, followed by characterization of the rewrite overheads in computational and financial terms.
-We finally discuss the impact of alternative platforms wrt database schema, database engine and LLMs.
+We finally discuss the impact of alternative platforms wrt database engine, database schema and LLMs.
 Rewrite Baselines.
@@
 Given an input query, each of the SOTA approaches is independently invoked to perform a rewrite, and the rewrite with the best performance is used as the baseline for comparison.
-Note that these approaches may generate rewrites that regress the performance [19].
-We discard such backsliding rewrites using, similar to LITHE, the Query-Costing module of the database engine.
-Query Set. Our evaluation uses the FPR subsets from the following diverse set of benchmarks, amounting to 100+ queries:
-  1. 1.
-    TPC-DS [11]: A synthetic OLAP benchmark modeling retail supplier environment with complex analytical queries.
-  2. 2.
-    DSB [12]: A TPC-DS variant with complex data distributions and additional query templates featuring many-to-many joins and non-equi-joins.
-  3. 3.
-    ARCHER [37]: A Text-to-SQL benchmark spanning 10 databases with availability of ground-truth SQL queries.
-  4. 4.
-    JOB [16]: An optimizer stress-test benchmark featuring queries with large and complex join graphs.
-  5. 5.
-    StackOverflow [20]: A real-world benchmark with query templates modeling questions and answers from experts. A random instance of each template is taken.
+Note that these approaches may occasionally generate rewrites that are expected by the optimizer’s costing module to regress the performance.
+For safety, we immediately discard such rewrites, similar to LITHE.
+Query Set.
+Our evaluation in this paper primarily focuses on complex analytical queries from the standard TPC-DS decision-support benchmark [12], which models a retail supplier environment.
+The benchmark is used at its default size of 100 GB.
+As mentioned earlier, we focus on “slow queries” that take over 10 seconds to execute in their original form, operating in a cold-cache environment.
+LITHE has also been evaluated on other benchmarks, including DSB [13], ARCHER [44], JOB [19] and StackOverflow [27]. Due to space limitations, we defer their details to Section A.
 Testbed.
-The majority of our experiments are carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA.
-Variations on this platform are considered in Section 7.4.
+The majority of our experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA.
+Variations on this platform are considered in Sections 7.4 and  7.5.
 Metrics.
-For each benchmark and each rewrite technique, we identified the number of queries for which a PR (productive rewrite) could be constructed.
-Subsequently, we computed the GM of the speedups obtained by each technique over the set of FPR (feasible productive rewrites).
-Finally, recall that if a rewrite achieves the maximum known speedup for the query, it is referred to as an MPR (Maximum Productive Rewrite).
-From the investment perspective, we measured the average rewrite time per query, and for the LLM-based techniques, the number of tokens used in the rewrite process.
-7.1 Rewrite Quality
-Table 6: Comparing LITHE with SOTA on FPR queries
-Benchmark (FPR) # PR SpeedupGM
-LITHE SOTA LITHE SOTA
-TPC-DS (65) 65 16 30.6 2.1
-DSB (20) 20 3 22.6 1.3
-ARCHER (25) 25 19 2.2 1.8
-JOB (5) 5 2 1.8 1.3
-StackOverflow (2) 2 1 9.0 7.5
-7.1.1 Benchmarks
-The speedup delivered by LITHE and SOTA on the FPR queries of the various benchmarks are shown in Table 6.
-The good news is that for TPC-DS and its DSB variant, LITHE produces PR for all the FPR queries – 65 in TPC-DS and 20 in DSB.
-In fact, most of these rewrites are highly productive as the SpeedupGM exceeds 30 for TPC-DS and 20 for DSB!
-Further, the PR coverage is far superior in comparison to SOTA– by a factor of 4 (TPC-DS) and 6 (DSB).
-Finally, the SpeedupGM of LITHE is higher by more than an order-of-magnitude wrt SOTA.
-Figure 7: Rewritten TPC-DS Q88 produced by LITHE
-7.1.2 Query Readability
-We now turn our attention to assessing how the readability of the queries is impacted by rewrites.
-That is, we wish to test whether the common wisdom that “efficient queries are more readable” holds for our rewrites.
-For this purpose, the number of joins in the query is used as a rudimentary metric to quantify query readability.
-The intuition is that a query with more joins is harder for the user to parse and understand.
-Table 7 presents the average number of joins in the original and rewritten queries for both SOTA and LITHE.
-The results clearly demonstrate that LITHE significantly reduces the number of joins, especially for TPC-DS, DSB and ARCHER benchmark, thereby enhancing query readability.
-Notably, for the DSB benchmark, SOTA actually worsened the readability, while LITHE maintained superior readability compared to the original query.
-However, for the JOB and StackOverflow benchmarks, where optimization opportunities were limited, the readability remained unchanged.
-Importantly, LITHE never performs worse than the original or SOTA-rewritten queries in terms of readability, always maintaining or improving the metric.
-Figure 7 shows the rewritten query for Q88 in which the number of joins reduced from 24 to 3.
-Overall, LITHE proves to be a reliable and effective tool for query rewriting, delivering both better performance and improved readability.
-Table 7: LITHE vs SOTA readability comparison
-Benchmark (FPR) Average # JOINS
-LITHE SOTA Original
-TPC-DS (65) 4.1 4.4 4.4
-DSB (20) 4.8 5.3 4.95
-ARCHER (25) 0.8 1 1.44
-JOB (5) 7 7 7
-StackOverflow (2) 8 8 8
-Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of FPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting.
-Nevertheless, LITHE continues to achieve PR in all of this limited set, whereas SOTA misses quite a few opportunities.
-Further, the SpeedupGM of LITHE is visibly better than SOTA.
-Figure 8: Speedups achieved by LITHE and SOTA
-Figure 9: Rewritten TPC-DS Q90
-Figure 10: Query execution times on TPC-DS 100GB
-7.1.3 Individual Queries
-The above results were for entire benchmarks. We now drill down into the performance at the granularity of individual queries. Due to space limitations, we focus only on the TPC-DS benchmark here.
-Figure 8 compares the performance of LITHE and SOTA on each of the 65 FPR queries in TPC-DS – note that the speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale.
-We first observe that, gratifyingly, rewrites are indeed capable of providing dramatic speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE.
+For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed.
+Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e.
+CPRs arising from either LITHE or SOTA).
+From the investment perspective, we measured the average rewrite time per query, and additionally for the LLM-based techniques, the number of tokens used in the rewrite process.
+7.1 Rewrite Quality (Cost and Time)
+7.1.1 Estimated Cost
+LITHE produces a rewrite with a positive speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold.
+Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5.
+On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1.
+All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27.
+We were able to prove the rewriting correctness of 9 CPRs through the logic-based QED tool.
+The remaining passed our result equivalence tests on both the down-sampled and full databases.
+Furthermore, as a final check, we manually verified the correctness of the rewritten queries.
+A drill-down into the cost speedup performance at the granularity of individual queries is shown in Figure 8, which compares LITHE (orange bars) and SOTA (blue bars) on each of the 27 CPR queries – note that the cost speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale, and the queries are sequenced in decreasing order of LITHE speedup.
+The vertical dotted line at $1$ represents the normalized baseline cost of the original query with the native optimizer, while the vertical line at $1.5$ is the CPR threshold.
+We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE.
 This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
-As another example, Figure 10 shows a rewrite where only LITHE achieves 2X speedup.
-Here, the two subqueries in the original query (each with 3 joins) are replaced by a common set of joins, and the aggregations are conditionally computed using CASE statements.
-Second, even among the 16 PR queries of SOTA, there are instances (e.g.
-Q6, Q47) where LITHE delivers substantially faster speedups than SOTA.
-Overall, across these 16 queries, LITHE attains a GM of 28.7 as compared to SOTA’s GM of 19.5.
-Finally, note that LITHE matched or improved on SOTA for all the FPR queries.
-7.1.4 Query Execution Times.
-Thus far, we had considered optimizer estimated execution costs.
-We now move to wall-clock runtimes for query executions – Figure 10 shows this metric (on a $\mathbf{log_{10}}$ scale) for the original query, LITHE, and SOTA on a representative set of queries executed on the TPC-DS 100GB database.
-We observe that substantial benefits are achieved, even exceeding order-of-magnitude improvements in some cases.
-Of course, as expected, these speedups may not directly match the optimizer-estimates because of the inherent imprecision in optimizer models.
-For example, the 5 orders-of-magnitude speedup estimated for Q41 decreases to a still respectable 3 orders-of-magnitude.
-The opposite also occurs: the 2 orders-of-magnitude speedup estimated for Q45 now becomes a huge 4 orders-of-magnitude.
-7.2 Ablation Analysis
+Figure 7: Plan Cost Speedups
+Figure 8: Execution Time Speedups
+Second, in most queries, the cost speedup of LITHE either exceeds or matches SOTA.
+In fact, there are several queries (e.g.
+Q45, Q25, Q4) where LITHE produces a highly beneficial CPR but SOTA simply returns the original query.
+Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query.
+And in Q88 and Q95, SOTA is marginally better than LITHE.
+At this stage, one could imagine that more rules could be added to LITHE to bring it on par with SOTA for queries like Q57.
+However, as discussed later in Section 8.3, such customized rules, while attractive at compile-time, run a significant risk of causing regressions at run-time.
+We have therefore selected only broad-brush rules for inclusion in LITHE.
+This conservative approach is justified below in the timing section, where we find that LITHE actually performed better than SOTA on these queries (Q57, Q88, Q95) wrt runtimes.
+7.1.2 Execution Time
+Thus far, we had considered optimizer-estimated execution costs.
+We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA.
+We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 10000!
+Second, in almost all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
+There remain only a few instances (Q9, Q47) where SOTA is marginally better.
+From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well.
+On the one hand, there is Q45 where the projected speedup of $10^{2}$ increases to a huge $10^{4}$ at runtime.
+On the other hand, the $10^{5}$ speedup for Q41 decreases to $10^{3}$.
+But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100.
+However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites.
+Overall, LITHE produces more robust rewrites resulting in a high TSGM of 18.4, whereas SOTA provides a TSGM of 6.
+7.1.3 Reasoning
+As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries.
+For example, consider Q45 – the provided explanation correctly lists the main reasons for the huge $10^{4}$ speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc.
+Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
+7.2 Ablation Analysis of Rewrite Quality
 7.2.1 Components of LITHE
 A natural question at this stage is the role of the various techniques in LITHE towards achieving its large performance benefits.
-This analysis is captured in Table 8 which lists the PR and MPR for each technique when invoked in isolation.
-The table also shows the cumulative number of PR and MPR when the different techniques are combined in the order of listing.
-Table 8: Ablation Analysis
+This analysis is captured in Table 5 which lists the CPRs for each technique when invoked in isolation as well as the cumulative number of CPRs when the different techniques are combined in the order of listing.
+Table 5: CPRs contributed by LITHE components.
+# CPR
 Isolated Cumulative
-# PR # MPR # PR # MPR
-Basic Prompts (Section 3) 19 13 19 13
-Rules R1 — R4 (Section 4.1) 17 12 23 15
-Rules R5, R6 (Section 4.2) 44 34 60 48
-MCTS (Section 5) 65 65 65 65
-We observe in the table that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) are capable of achieving maximum speedups for around a quarter (15/65) of the FPR query set.
-When the statistics-infused Rules (R5, R6) are added, this fraction jumps to about three-quarters (48/65).
-And finally, when MCTS is included, we reach full coverage of the FPR.
-Table 9: Rewrite Overheads of LITHE and SOTA on FPR queries
-Benchmark Avg. Time (min) Avg. Tokens Avg. Cost USD
-LITHE SOTA LITHE SOTA LITHE SOTA
-TPC-DS 4.6 1.7 22390 20076 0.055 0.050
-DSB 8.7 4.0 22203 15699 0.055 0.039
-ARCHER 2.3 0.6 7284 5465 0.018 0.013
-JOB 6.1 1.3 22958 13692 0.057 0.034
-StackOverflow 7.1 3.2 22634 12759 0.056 0.032
+Basic Prompts (Section 3) 15 15
+Rules R1 — R4 (Section 4.1) 10 18
+Rules R5, R6 (Section 4.2) 11 25
+MCTS (Section 5) 26 26
+We observe that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) contribute to around two-thirds of the CPRs (18/26).
+When the statistics-infused Rules (R5, R6) are added, this number jumps to 25.
+Finally, MCTS produces an additional CPR and also improves the cost speedup of a pre-existing CPR.
 7.2.2 Database-Sensitive Prompts.
@@
 The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads.
-When this experiment was conducted, the SpeedupGM dropped precipitously to a paltry 2.3, a far cry from the 30.6 obtained with the database-sensitive rules.
+When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules.
 These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
-7.3 Rewrite Overheads
-Having established the ability of LITHE to deliver performance-beneficial rewrites, we now turn our attention to the overheads incurred in this rewriting process.
-Here, in addition to rewrite time, the monetary outlay for LLM inferencing also has to be considered.
+7.3 Rewrite Overheads (Time/Money)
+Having established the performance benefits of rewrites, we now turn our attention to their time and financial overheads.
 7.3.1 Transformation Time
-The average processing time per FPR query is shown in Table 9 for LITHE and SOTA over the various benchmarks.
-The good news is that the rewrites are typically identified in a few minutes even for highly complex queries.
-We hasten to note that this investment is likely to be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads.
-As a case in point, for TPC-DS Q30 on a 10GB database, the original version took 17 minutes to run to completion, whereas the LITHE rewrite took 7 minutes to transform but finished in less than a second!
-Further, when the database size was increased to 100GB, the LITHE rewrite executed in just 8 seconds, while the original version took more than an hour.
-Finally, many applications tend to use a set of canned queries which are run thousands of times.
-Thus, a relatively small one-time investment can be easily recovered over repeat executions of such queries.
+The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes.
+However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads.
+For instance, with Q11, the original query took nearly two hours to complete, whereas the LITHE rewrite executed in just 3 minutes.
+Further, many applications tend to use a set of canned queries which are run thousands of times.
+Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
+Table 6: Rewrite Overheads of LITHE and SOTA.
+Avg. Time (min) Avg. Tokens Avg. Cost (USD)
+LITHE 5 18427 0.045
+SOTA 1.7 20076 0.050
 Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites.
-A drill-down into the various components of LITHE is shown in Table 10 for the TPC-DS queries.
-We observe here that the lion’s-share is taken by the initial prompt ensemble and the final MCTS module, largely arising from their “brute-force” combinatorial construction.
-But both these modules can potentially be improved – the choice of prompt via a classifier, and MCTS by early pruning of ill-formed (semantic/syntactic errors) queries.
-These options are discussed below in Section 7.3.3.
-Table 10: LITHE Transformation Time Analysis
-Time/Query (min) Tokens / Query
-Avg. Max. Avg. Max.
-Basic Prompts (Section 3) 1.2 5.2 2407 17778
-Rules R1 — R4 (Section 4.1) 0.8 3.6 3076 11988
-Rules R5, R6 (Section 4.2) 0.5 2.1 2114 5865
-MCTS (Section 5) 2.1 10.6 14792 55055
+A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module.
+We explore techniques to improve their efficiency in the Section B.
 7.3.2 Monetary Outlay
-The average number of LLM tokens required by LITHE and SOTA for the various benchmarks, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are shown in Table 9.
+The average number of LLM tokens required by LITHE and SOTA, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are also shown in Table 6.
 The good news is that the inference charges per query are just a few cents, making rewriting practical from a deployment perspective.
-We also see that LITHE can be a factor of two more expensive than SOTA.
-To understand the source of this expense, the token costs incurred by each module of LITHE are shown in Table 10.
-Again, the lion’s share is taken by the MCTS module.
-This is to be expected due to its exploratory nature which can result in traversal of several paths in the tree – however, the additional expense appears to be justified by its output of significantly better rewrites.
-Finally, even this expense gets reduced when the time overheads are cut down, as described below.
-7.3.3 Overheads Reduction
-Reducing the query rewrite time while still obtaining the same performance would be possible if we could directly use the MCTS-driven rewrite with the appropriate prompt.
-Towards this end, we build a classifier to pick the most appropriate rewrite rule for a given input query.
-Specifically, the classifier identifies which, if any, of Rules R1–R6 is appropriate for a given query.
-If none are appropriate, then it falls back to just the set of basic prompts to identify the best prompt to be given as input to the MCTS module.
-In addition to reducing the rewrite times, using a classifier can also reduce the financial costs of the rewrite.
-We design an LLM based classifier to accomplish this as follows.
-The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied.
-For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision.
-Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
-Table 11 compares the performance of LITHE with and without the classifier.
-The time overheads do visibly go down by about 40 percent, and the tokens by about 20 percent.
-However, there is a price to be paid – the PR is reduced to 59 and the SpeedupGM comes down to 21.7.
-In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
-Pruning in MCTS.
-A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output.
-In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early.
-We are currently working on the design and implementation of such checks.
-7.4 Alternative Platforms (DB/Engine/LLM/T2S)
-7.4.1 Anonymized Database
-An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of the GPT-4o LLM having already been trained well on our evaluated benchmarks, all of which are in the public domain.
-To investigate this issue, we created an anonymized version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their content.
-We then constructed rewrites for the 65 FPR queries (after syntactic changes to reflect the new schema) on this version.
-The results are shown in Table 12 for SOTA and LITHE.
-We observe that while the SpeedupGM marginally decreases, the performance profiles remain similar to those seen earlier for native TPC-DS, testifying to the robustness of the rewrite approaches.
-Table 11: Impact of Classifier (TPC-DS)
-Metrics No Classifier With Classifier
-# PR 65 59
-SpeedupGM 30.6 21.7
-Avg. Tokens 22390 18296
-Avg. Time 4.6 min 2.8 min
-Table 12: Rewrite Performance on Anonymized Database
-Approach # PR SpeedupGM
+7.4 Commercial DBMS
+A legitimate question could be whether the rewrites made amends for the PostgreSQL optimizer but may fail to be useful in highly-engineered database engines.
+To evaluate this issue, we performed TPC-DS rewrites on a pair of popular commercial DBMS, which we refer to as OptA and OptB.
+Table 7: Performance on Commercial Database Engines.
+# CPR CSGM TSGM
+OptA OptB OptA OptB OptA OptB
+LITHE 7 8 3.6 4.8 1.5 2.4
+SOTA 3 6 2.1 3.5 1.3 1.7
+The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA.
+Despite the apparent lack of optimization headroom, LITHE still produces 7 and 8 CPRs resulting in a healthy CSGM of 3.6 and 4.8, respectively.
+Further, the TSGM provided by these rewrites are a useful 1.5 and 2.4, respectively.
+However, unlike the no-regression outcome seen with PostgreSQL, LITHE generates 3 CPRs on OptA, and 2 on OptB, which turn out to be regressions.
+So, although these commercial optimizers may have greater sophistication, they also suffer significant brittleness with regard to estimation accuracy.
+Interestingly, one regression on OptB was caused by the LLM bringing in, via a basic prompt, an unrestricted version of Rule 6, which is prone to generating poorly-estimated plans (see Section 8.3).
+Despite such occasional errors, LITHE provides good overall improvement in runtimes as indicated by the TSGM values.
+We also performed a preliminary study of LITHE on an internal real-world benchmark, whose data and query characteristics are more realistic than TPC-DS.
+Even on this hard benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with a CSGM of 2 and a TSGM of 3.3.
+The above results suggest LITHE has a useful role to play in industrial environments.
+From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
+7.5 Dependence on LLM (Training/Model)
+7.5.1 Masked Database
+An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain.
+To investigate this issue, we created a masked version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their contents.
+We then constructed rewrites for the CPR queries (after syntactic changes to reflect the new masked schema) on this version.
+The results are shown in Table 8, and we observe that the performance profiles only marginally decrease for both LITHE and SOTA, testifying to their robust applicability.
+Table 8: Rewrite Performance on Masked Database.
+Approach # CPR CSGM
 TPC-DS Masked TPC-DS Masked
-SOTA 16 14 2.1 1.8
-LITHE 65 64 30.6 29.4
-7.4.2 Commercial DBMS
-Another legitimate question could be whether the rewrites made amends for the relative lack of sophistication in the PostgreSQL optimizer, and that its benefits would disappear in premier database engines.
-To evaluate this issue, we performed TPC-DS rewrites on two popular and highly engineered commercial DBMS.
-Table 13 compares the performance of LITHE and SOTA on these systems.
-Compared to PostgreSQL, the PR reduced considerably from 65 to 15 and 29, respectively, in case of LITHE, while SOTA had just 6 and 14 PRs respectively.
-Nevertheless, a significant fraction of the benchmark queries are still represented, providing a very respectable SpeedupGM of 5.3 and 3.4.
-So, LITHE has a role to play even on high-end database engines.
-From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
-Table 13: Rewrite performance on Commercial Database Engines
-Approach # PR SpeedupGM
-Opt A Opt B Opt A Opt B
-LITHE 15 29 5.3 3.4
-SOTA 6 14 2.2 2.5
+SOTA 15 12 6.1 5.3
+LITHE 26 24 11.5 10.5
 7.5.2 LITHE on LLaMA
-In our concluding experiment, we evaluate the performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which is said to have several hundred billion parameters [5].
+In our concluding experiment, we evaluated LITHE’s performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which may have several hundred billion parameters [5].
 To attain practical inference times, the model was loaded using a low 4-bit quantization.
 Further, to ensure reproducibility and deterministic answers, the do_sample parameter was set to False, which forces the LLM to perform greedy decoding.
-In order to make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
-For this environment, Table 14 shows the PR and SpeedupGM obtained on the micro-benchmark with and without MCTS.
-Although certainly lower than the corresponding numbers with GPT-4o (see Tables 2 and  5), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support.
-So, the message is that even smaller models can be fruitfully used in real-world environments.
-Table 14: Micro-benchmark Performance with LLaMA
-# PR SpeedupGM
-LLaMA without MCTS 5 2.6
-LLaMA with MCTS 9 7.2
-7.4.4 Text-to-SQL
-As mentioned in the Introduction, we could alternatively construct the SQL-to-SQL rewrite problem as a two-stage pipeline: SQL-to-Text followed by Text-to-SQL.
-To assess this option, we first implemented SQL-to-Text on GPT-4o with the prompt asking for conversion of the SQL query to a clear business description.
-Then, two open tools were evaluated for Text-to-SQL – CHESS [28] and SQLCoder [3], which have performed very well on the BIRD [17] and SQL-Eval [2] benchmarks, respectively.
-However, in all our experiments, the outcomes were uniformly poor with the rewrites often suffering semantic errors, and even among the few correct rewrites, there was not a single PR.
-Therefore, this alternative architecture does not appear to be a viable option.
+To make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
+For this environment, Table 9 shows LITHE’s performance on the TPC-DS benchmark with and without MCTS.
+Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support.
+So, the message is that smaller models can also be fruitfully used in real-world environments.
+Table 9: LITHE Rewrite Performance with LLaMA.
+# CPR CSGM TSGM
+LLaMA without MCTS 17 5.6 8.2
+LLaMA with MCTS 21 8.3 14.7
+8 The Road Ahead
+Based on our study, we now present a few observations with implications for the future design and deployment of rewriting tools.
+8.1 Rewrite Space Coverage by LLMs
+Given the decades-long research on database query optimization, we expected the potential for performance improvement via rewriting to be limited.
+What came as a surprise was the substantial scope for improvement still available, as showcased by the large CSGM and TSGM values, even on commercial platforms.
+These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines.
+Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [7].
+8.2 Rewrite Migration to Optimizer
+The above demonstrated the potent exploratory power of LLMs.
+But from an overheads perspective, such rewrites should ideally be within the optimizer’s native search space rather than recommended from outside.
+Therefore, it would be a useful exercise to try and distill fresh optimization rules from these instances, leveraging the extensibility features of contemporary optimizers [14] to facilitate their incorporation in existing systems.
+Figure 9: Rewritten TPC-DS Q90
+On the flip side, there appears to be an “impedance mismatch” against such integration for certain classes of rewrites.
+For example, consider the TPC-DS Q90 rewrite in Figure 9.
+The original query individually computed AM (morning) sales and PM (evening) sales, which were then used to compute the AM to PM ratio.
+The rewrite, however, extracted all relevant rows in one shot and computed the ratio using CASE statements – encoding such transformations as generic rules in the optimizer appears challenging, given the combinatorial ways in which such transformations can occur.
+Therefore, a fruitful area of future research could be achieving a middle-ground between the disparate world-views of LLMs and traditional optimizers.
+8.3 Revisiting Optimizer Plan Costing
+As highlighted in Section 7, there were instances of substantive differences between the promised speedup and that delivered at run-time.
+In fact, to the extent that speedups could even turn out to be regressions!
+This is due to the brittleness of optimizer plan costing in the new spaces explored by the LLM.
+As a case in point, during our LITHE design process, the prefiltering in Rule R6 (Table 1) had initially not been restricted to self-joins.
+It resulted in the number of CPRs (on PostgreSQL) being as high as 65, with an astonishing CSGM of 30.6!
+However, upon execution, most of these rewrites turned out to be regressions, which led to our inclusion of the restriction.
+But note that we are artificially modulating the rewrite, rather than fixing the plan costing module, which is the principled solution.
+In sum, while plan cost modeling has been a long-standing area of research, there is now even more reason given the new rewrite spaces to study this topic further – for instance, the operator cost model could be extended using calibration techniques similar to those advocated in [40], while the operator cardinality model could be improved with attention-based techniques [23].
+8.4 Agentic LLMs for Query Rewriting
+An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment.
+This would allow a “LITHE agent” to leverage a memory store to log one or more of: (1) query execution times; (2) memory usages of slow queries; (3) prior interactions with users; and (4) useful/new rewrite rules learnt over time.
+Additionally, the agent could be given database engine access so that it can directly request metadata such as query execution plans, database statistics, live query analyzers, etc.
+By following a structured workflow – planning, acting, observing, and refining – the agent would then iteratively enhance query performance.
+This adaptive workflow supports continuous learning, so that query rewriting can become more efficient over time.
+8.5 Scope of Semantic Equivalence Tools
+As seen in the experiments section, the coverage of logic-based query equivalence testing on industrial-strength queries is rather limited.
+On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness.
+This limitation restricts the use of LITHE in a fully automated scenario, i.e., as a direct preprocessor to the query engine.
+Therefore, a key challenge is to improve logic-based coverage.
 9 Related Work
@@
 R-Bot [32] also leverages an LLM to optimize the order of Calcite rules, but employs advanced contemporary techniques such as retrieval-augmented generation (RAG) and step-by-step self-reflection to improve the outcomes.
-Query Booster [7] implements human centred rewriting – it provides an interface that the user can use to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query.
-There also exist other rule-based rewrite approaches that are designed for specific type of rewrites such as optimizing correlated window aggregations [32] and common expression elimination [9].
+Query Booster [8] implements human-centered rewriting – it provides an interface to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query.
+There are also rule-based rewrite approaches designed for specific types of rewrites such as optimizing correlated window aggregations [39] and common expression elimination [10].
 All of the above approaches operate via the query plan space, which can restrict the kind of rewrites that can be accomplished.
@@
 A limitation, however, is that LLM-generated rewrite rules often fail to generalize beyond specific query pairs.
-Even when generalized rules are present, without accompanying examples, LLMs can struggle to apply rules correctly.
+Even when generalized rules are present, LLMs can struggle to apply rules correctly if not provided with accompanying examples.
 Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
-LLMs for Text-to-SQL.
-LLMs have also been extensively used for Text-to-SQL transformations [17, 33, 2, 3, 28].
+LLMs for Database Modules.
+LLM technologies have been advocated for a variety of database modules.
+For instance, they have been extensively used for Text-to-SQL transformations [22, 41, 2, 3, 33, 46].
 The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [35, 29, 33].
-On the other hand, the goal of SQL rewriting is on improving the performance of an existing SQL query.
+On the other hand, the goal of S2S rewriting is on improving the performance of an existing SQL query.
 Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
+In recent times, LLMs have also been considered for plan-hinting [7], join-order optimization [34], index selection [46], data pipelines [18], data management [21], etc.
+An LLM-enabled multi-modal query optimizer [37], where data spans text, image, audio and video domains, has also been proposed recently.
+These approaches can be used in conjunction with the LITHE system since they address orthogonal segments of the query processing pipeline.
 10 Conclusions and Future Work
@@
 An empirical evaluation over common database benchmarks showed that rewriting is a potent mechanism to improve query performance.
-The speedups obtained are significant and proof-of-concept, but there remains scope for improvement.
-For instance, while a speedup of 83X was produced for Q95 of TPC-DS, we were able to hand-craft an alternate rewrite with 99X speedup by moving a self-join from a CTE to the main query, thereby reducing scans of the web_sales table.
-We are currently working on adding a few select rules to implement such optimizations within the LITHE framework.
-Our focus here was primarily on prompting-based strategies.
-In our future research, we also plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
+In fact, even order-of-magnitude speedups were routinely achieved with regard to both abstract costing and execution times.
+However, our results also showed a significant semantic distance between foundation models and query optimizers, with regard to both scope and precision, which would have to be bridged to fully leverage the latent power of LLMs.
+Further, our focus here was primarily on prompting-based strategies – in our future research, we plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
 Appendix
-A1 : Detailed MCTS Algorithm
+Appendix A Rewrite Quality (Other Benchmarks)
+We also evaluated LITHE on the following set of benchmarks.
+  1. 1.
+    DSB [13]: A TPC-DS variant with complex data distributions and additional query templates featuring many-to-many joins and non-equi-joins.
+  2. 2.
+    ARCHER [44]: A Text-to-SQL benchmark spanning 10 databases with availability of ground-truth SQL queries.
+  3. 3.
+    JOB [20]: An optimizer stress-test benchmark featuring queries with large and complex join graphs.
+  4. 4.
+    StackOverflow [27]: A real-world benchmark with query templates modeling questions and answers from experts. A random instance of each template is taken.
+Note that the queries in these workloads (with the exception of DSB) are mostly fast running (i.e.
+completing in less than 10 seconds).
+Therefore, with the primary goal to test the coverage of LITHE over a diverse variety of workloads (and not the deployment angle), we include all the queries in our analysis.
+The number of CPRs produced and the cost speedup delivered by LITHE and SOTA over these benchmarks are shown in Table 10.
+Table 10: Comparing LITHE with SOTA on different benchmarks.
+Benchmark # CPR CSGM
+LITHE SOTA Union LITHE SOTA
+DSB 9 3 9 7.7 1.7
+ARCHER 22 19 22 2.1 1.95
+JOB 4 2 4 1.9 1.4
+StackOverflow 2 1 2 8.7 7.5
+In case of DSB, LITHE produces CPR for 9 queries resulting in a highly productive CSGM of 7.7.
+Similar to the case with TPC-DS, LITHE performs better than SOTA both with respect to its coverage, as well as the cost speedups.
+Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of CPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting.
+Nevertheless, LITHE continues to achieve better CPR coverage, whereas SOTA misses quite a few opportunities.
+Further, the CSGM of LITHE is visibly better than SOTA.
+Appendix B Overheads Reduction for LITHE
+Rule Selection Classifier
+Reducing the query rewrite time while still obtaining the same performance would be possible if we could directly use the MCTS-driven rewrite with the appropriate prompt.
+Towards this end, we build a classifier to pick the most appropriate rewrite rule for a given input query.
+Specifically, the classifier identifies which, if any, of Rules R1–R6 is appropriate for a given query.
+If none are appropriate, then it falls back to just the set of basic prompts to identify the best prompt to be given as input to the MCTS module.
+In addition to reducing the rewrite times, using a classifier can also reduce the financial costs of the rewrite.
+We design an LLM based classifier to accomplish this as follows: The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied.
+For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision.
+Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
+Table 11: Impact of Classifier (TPC-DS)
+Metrics Without Classifier With Classifier
+# CPR 26 23
+CSGM 11.5 8.5
+Avg. Tokens 18427 16003
+Avg. Time 5 min 2.4 min
+Table 11 compares the performance of LITHE with and without the classifier.
+The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent.
+However, there is a price to be paid – the CPR is reduced to 24 and the CSGM comes down to 8.5.
+In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
+Pruning in MCTS
+A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output.
+In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early.
+We are currently working on the design and implementation of such checks.
+Appendix C Query Equivalence Testing via Sampling
+If the original query is not within the scope of the logic-based QED tool, we alternatively use a sampling-based approach to test equivalence.
+The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results.
+However, while this test is a necessary condition for query equivalence, it is not a sufficient condition.
+That is, there are no false negatives, but there can be false positives.
+This is because the sampled database may not cover all the predicates present in the query.
+This can cause two types of problems: First, it is possible for two different queries to return the same result.
+This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other.
+Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries.
+This again does not imply that the queries are equivalent.
+To statistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples.
+The goal is to reduce the likelihood of non-equivalent queries returning the same results.
+To minimize the occurrence of the second problem (empty results), the following approach is taken:
+  1. 1.
+    We use correlated sampling [42] to sample the database.
+This technique leverages the join graph of the schema to produce a sample that maintains join integrity between the tables participating in the query.
+  2. 2.
+    To differentiate between inner and outer joins, we insert into the sample database, rows with appropriate values on the relevant columns such that they produce NULL values for outer joins.
+  3. 3.
+    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result.
+We make use of the rows in the sampled data for this purpose.
+For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database.
+We then replace the query constant with a value already present in the sample.
+Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database.
+Note that these modifications only change the selectivity of the query, but not its semantics.
+Appendix D Detailed MCTS Algorithm
 Algorithm 2 Token-augmented Rewrite
@@
 45:    return the original query
-A2 : Query Equivalence Testing
-We use a multi-stage approach, described below, to test semantic equivalence between the original query and a candidate rewrite.
-1.
-Logic-based Equivalence.
-Although verifying the equivalence of a general pair of SQL queries is a long-standing hard problem [6], a variety of logic-based tools are available for proving equivalence over restricted classes of queries (e.g.
-Cosette[10], SQL-Solver [13], QED [30]).
-In our pipeline, we used the recently proposed QED [30] since it was found to cover a larger set of queries compared to the alternatives.
-The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
-2.
-Result Equivalence via Sampling.
-If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence.
-The idea here is to execute the queries on a small sample of the database and verify equivalence based on the sample results.
-However, while this test is a necessary condition for query equivalence, it is not a sufficient condition.
-That is, there are no false negatives, but there can be false positives.
-This is because the sampled database may not cover all the predicates present in the query.
-This can cause two types of problems: First, it is possible for two different queries to return the same result.
-This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other.
-Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries.
-This again does not imply that the queries are equivalent.
-To probabilistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples.
-The goal is to reduce the likelihood of non-equivalent queries returning the same results.
-To minimize the occurrence of the second problem (empty results), the following approach is taken:
-  1. 1.
-    We use correlated sampling [34] to sample the database.
-This technique leverages the join graph of the schema to produce a sample that maintains join integrity between the tables participating in the query.
-  2. 2.
-    To differentiate between inner and outer joins, we insert into the sample database, rows with appropriate values on the relevant columns such that they produce NULL values for outer joins.
-  3. 3.
-    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result.
-We make use of the rows in the sampled data for this purpose.
-For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database.
-We then replace the query constant with a value already present in the sample.
-Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database.
-Note that these modifications only change the selectivity of the query, but not its semantics.
-3.
-Result Equivalence on the Entire Database.
-Instead of resorting to samples, result equivalence could also be evaluated, in principle, on the entire database itself.
-However, this could turn out to be prohibitively expensive, especially if the queries themselves are time consuming (e.g., due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions.
-Therefore, this option is disabled by default in our workflow, but the user can optionally invoke it as a final verification step after the sampled equivalence testing.
-A3 : Examples used in Rules 1–6
+Appendix E Examples used in Prompts for Rules 1–6
 R1: Use CTEs (Common Table Expressions) to avoid repeated computation.
@@
 order by item.id;
-R6: Pre-filter fact tables in a CTE using those dimension tables having low selectivity values. Retain dimension table filters in main query. Do not create explicit join statements.
+R6: Pre-filter tables that are involved in self-joins and have low selectivities on their filter and/or join predicates.
+Remove any redundant filters from the main query.
+Do not create explicit join statements.
 Original Query
-Select item.id, item.code, item.price, country.name
-from item, element, country, colour
-where element.sourceid = item.sourceid
-  and item.code = country.code
-  and item.colorcode = colour.colorcode
-  and element.zip > 8000
-  and country.name = ’US’
-  and colour.colour_name = ’Red’
-order by item.id;
+with total_price_cte as (
+    select item.id, colour.colorcode, sum(item.price) total_price
+    from item, color
+    where item.colorcode = colour.colorcode
+    group by item.id, colour.colorcode
+select t_sec.id, t_first.colorcode
+from total_price_cte t_first, total_price_cte t_sec
+where t_sec.id = t_first.id
+    and t_first.colorcode = ‘R’
+    and t_sec.colorcode = ‘R’
+    and t_first.total_price > 0
+order by t_sec.id
+limit 100;
 Statistics
 Selectivity of different predicates is given below :
-( 1 ) element.zip > 8000 :: 0.002
-( 2 ) country.name = ’US’ :: 0.4
-( 3 ) colour.colour_name = ’Red’ :: 0.01
+( 1 ) colour.colorcode = ‘R’ :: 0.01
 Rewritten Query
-with filtered_item as(
-  Select *
-  from item
-  where
-    item.sourceid in (
-      Select
-        element.sourceid
-      where
-        element.zip > 8000
-    and item.colorcode in (
-      Select
-        colour.colorcode
-      from
-        colour
-      where
-        colour.colour_name = ’Red’
-Select item.id, item.code, item.price, country.name
-from filtered_item as item, element, country, colour
-where
-  item.sourceid = element.sourceid
-  and item.code = country.code
-  and item.colorcode = colour.colorcode
-  and element.zip > 8000
-  and country.name = ’US’
-  and colour.colour_name = ’Red’
-order by item.id;
+with total_price_cte as (
+    select item.id, colour.colorcode, sum(item.price) total_price
+    from item, color
+    where item.colorcode = colour.colorcode
+    group by item.id, colour.colorcode
+filtered_total_price_cte as (
+    select * from total_price_cte
+    where colorcode = ‘R’
+select t_sec.id, t_first.colorcode
+from filtered_total_price_cte t_first, filtered_total_price_cte t_sec
+where t_sec.id = t_first.id
+    and t_first.total_price > 0
+order by t_sec.id
+limit 100;

### v3

@@
 The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer.
-For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
+For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
 Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise query performance.
@@
 This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete.
-However, the query can be equivalently rewritten (assuming NOT NULL column constraints) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
+However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
 For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives.
@@
 Prior Work
-A variety of innovative Rule-based (e.g.
+A range of innovative Rule-based (e.g.
 Learned Rewrite [48]) and Model-based (e.g.
-Gen-Rewrite [25]) approaches have been proposed in the literature for S2S transformations.
-Further, a recent vision paper [46] has persuasively advocated the use of LLMs for such query rewriting.
-These novel state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting.
-However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space – the latter provides greater scope for candidate rewrites since only intent, and not implementation, is expressed.
+Gen-Rewrite [28]) approaches have been proposed for S2S transformations.
+A recent vision paper [49] also advocates the use of LLMs for such query rewriting.
+The state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting.
+However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e transforming the query structure itself).
 In this paper, we address these lacuna and amplify the effectiveness of query rewriting.
+In particular, our query space-based transformations provide greater scope for candidate rewrites since only intent, and not implementation, is expressed, and this multiplicity of rewrites in turn results in a broader coverage of the underlying plan space.
 Figure 3: High-level architecture of LITHE
@@
 Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$.
+Note that having the DBA in the loop is a common practice in commercial query advisory systems [14].
 Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
@@
 Calcite [10] rules) typically used in existing rewrite systems, which operate in plan space.
-In particular, we work with two classes of rules:
-1. Redundancy Removal Rules: These rules are meant to eliminate repeated and redundant computations of the same output.
-2.
+In particular, we work with two classes of rules – 1.
+Redundancy Removal Rules: These rules eliminate repeated and redundant computations of the same output; and 2.
 Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts.
@@
 The above rewrite options are restricted to the standard prompt interface.
-Additionally, we leverage the rich telemetry provided by LLMs – in particular, the token probabilities output at each step in the prediction sequence.
+We also leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence.
 Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process.
 To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation.
-MCTS has previously demonstrated significant effectiveness in code generation exercises [43].
+MCTS has previously been used to improve ordering of query rewrite rules [48] in the plan space, but our usage is new – it is geared to guide the LLM in exploring rewrites in the query space.
 Semantic Equivalence Labels.
@@
 Unfortunately, the general problem of proving query equivalence is NP-complete [7].
-However, for restricted classes of queries, various logic-based equivalence-proving tools are available, and in LITHE, we use the recently developed QED tool [36] to verify rewrite correctness for queries within its scope.
-The remaining out-of-scope queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database.
-Finally, result equivalence can be optionally tested on the original database itself.
+We therefore take a twofold approach to check for query equivalence.
+First, the queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database.
 While these checks do not suffer from false negatives, they may (rarely) incur false positives.
-After carrying out the above exercise, LITHE provides a verification label – provable or statistical – regarding semantic equivalence.
+These tests can thus quickly weed out obviously non-equivalent rewrites.
+A final check is then performed using logic-based equivalence-proving tools that provably work for a restricted classes of queries.
+Specifically, LITHE uses the recently developed QED tool [40] and SQLSolver [17] to verify rewrite correctness for queries within their scope.
+Whereas, for out-of-scope queries, result equivalence on the original database itself is used to test the equivalence.
+After carrying out the above exercise, LITHE provides a verification label of semantic equivalence – provable if verified by a logic-based tool, or statistical otherwise.
 Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
+Regression Identification.
+Finally, it is well-known that in practice there are often material discrepancies between optimizer cost predictions and actual execution times [29].
+So it is possible that a promising rewrite as per the optimizer may actually regress at run-time as compared to the original formulation.
+To help guard against such possibilities, we have implemented heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
 Experimental Evaluation
@@
 Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
-We also evaluated whether the above cost reductions translated to real execution speedups.
+We also evaluated whether the above cost reductions translated into real execution speedups.
 Here, we find that LITHE is indeed often substantively faster at run-time as well.
-Specifically, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
-However, we also found that the well-known discrepancies between optimizer cost predictions and actual execution times [26] could be exacerbated by an injudicious choice of rule formulations.
-This separation suggests a future line of research to reduce the reasoning distance between language-based foundation models and statistics-based optimizer models.
+Specifically, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
+Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
 Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
@@
 3.
-Semantic Equivalence.
-A suite of provable (logic-based) and statistical (result-based) techniques are employed to assess the semantic equivalence of a recommended rewrite.
+Fast Semantic Equivalence.
+Statistical (result-based) techniques are employed to quickly and cheaply assess the semantic equivalence of a recommended rewrite.
 If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
@@
 5.
-Output.
-The least expensive valid rewrite as identified by the MCTS module is returned to the DBA together with the cost of the rewritten query and the approach used to claim equivalence.
-Additionally, an LLM-generated reasoning for the expected performance improvement is also provided.
-If no valid rewrite is identified, the original query itself is returned.
+Final Checks and Output.
+The least expensive valid rewrite (as identified by the MCTS module) is tested using a suite of provable (logic-based) techniques to generate a final equivalence label (provable or statistical), and its cost benefit over the original query is computed.
+Further, the execution “brittleness” is assessed using the robustness heuristics.
+If no valid rewrite is identified, or if the rewrite is expected to be a regression, the original query itself is returned to the DBA.
+Whereas, if a beneficial rewrite is recommended, an LLM-generated reasoning for the expected performance improvement is also extracted.
 2.1 Performance Framework
@@
 Performance
-The performance of the four prompt templates on the micro-benchmark is shown in Table 1.
+Table 1 shows the performance of the four prompt templates on the micro-benchmark.
 We find that less than half the rewrites are productive with individual prompts.
@@
 This approach would benefit queries with multiple types of redundancies.
-However, it also increases rewrite overheads due to repeated prompt applications.
+However, it also introduces significant computational overheads due to the vast number of possible rule application orders, making it expensive to explore.
 So, for simplicity, we have chosen to process them individually rather than cumulatively.
@@
   1. 1.
-    Selectivities of columns appearing in WHERE and JOIN clauses, as obtained from the database engine.
+    Estimated selectivities of columns appearing in the WHERE and JOIN clauses – these values are obtained via calls to the cardinality estimation modules of the query optimizer.
   2. 2.
@@
 We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3.
-Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in the Section B.
+Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in Section B.
 5 Token Probability Driven Rewrite
-A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up.
-This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs.
+A key challenge with LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up.
+This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs [20].
 In order to have a robust approach for such cases, we take inspiration from the code generation literature [46].
 Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
+As mentioned earlier, MCTS has been previously used for ordering rule applications in query rewrites [48], whereas our goal is to guide the LLM in exploring query rewrites.
 This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [33].
@@
 10:    while $len(n_{cur}.children)>0$ do
-11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$
+11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
 12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
@@
 Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$.
-This node is expanded in a greedy fashion, based on the highest-probability tokens until a terminal node is reached (Line 21).
+This node is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21).
 Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23).
-For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides with respect to the original query.
+For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides w.r.t the original query.
 However, if invalid (i.e.
@@
 After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
-4.
-Back Propagation: The $V$ value of the simulation for $n_{cur}$ is backpropagated to all its ancestor nodes.
-An ancestor node’s $V$ value is updated if and only if the new value is higher than the existing value.
+4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is back-propagated to all its ancestor nodes, and their values are updated iff the new value is higher than the existing value.
 Rewritten Query.
@@
 5.3 Performance
-We find that LITHE’s performance with MCTS-based rewrites fully matches the human target, providing 10 CPRs and the maximum CSGM of 11.84.
-While these improvements, as compared to just prompting, might seem marginal at this stage, we wish to make two observations: (1) MCTS helps extract the maximum improvement wherever possible; and (2) As shown later in Section 7.5.2, improvements due to MCTS become significant for smaller LLMs such as LLaMA.
+LITHE’s performance with MCTS-based rewrites fully matches the human target, yielding 10 CPRs and the maximum CSGM of 11.84.
+While these gains over just prompting may seem marginal, we wish make two observations: (1) MCTS extracts maximum improvement whererever possible, and (2) As shown in Section7.5.2, its impact becomes significant for smaller LLMs like LLaMA.
 6 Implementation Choices
@@
 6.1 LITHE Parameter Settings
-The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response.
-While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible.
-Hence we set the temperature to 0 which forces the model to greedily sample the next token.
+The “temperature” parameter of GPT-4o, which ranges over [0,1], controls randomness of the model’s response.
+While higher values suit creative tasks that require diverse and exploratory outputs, we seek focused, deterministic answers.
+Thus, we set the temperature to 0, forcing the model to greedily sample the next token.
 The hyperparameters used by LITHE for MCTS are as follows: The maximum number of iterations $iter_{max}$ is set to 8, expansion threshold $\theta$ is 0.7, and number of expansions $k$ is 2.
@@
 6.2 Query Equivalence Testing
-We use a multi-stage approach, described below, to test semantic equivalence between the original query and a candidate rewrite.
+We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite.
+Note that, as mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems[14].
 1.
-Logic-based Equivalence.
-Although verifying the equivalence of a general pair of SQL queries is NP-complete [6], a variety of logic-based tools (e.g.
-Cosette[11], SQL-Solver [15], VeriEQL [17], QED [36]) are available for proving equivalence over restricted classes of queries, as mentioned in the Introduction.
-In LITHE, we use the recently proposed QED [36] since it was found to cover a larger set of queries compared to the alternatives.
-The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
-2.
 Result Equivalence via Sampling.
-If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence.
+We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline.
 The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results.
@@
 That is, false positives may be present because the sampled database may not cover all the predicates featured in the query.
-To minimize this likelihood, we use a combination of (1) correlated sampling [42] for maintaining join integrity in the sample, (2) adding synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Section C.
+To minimize this likelihood, we use a combination of (1) correlated sampling [45] for maintaining join integrity in the sample, (2) adjusting constants in the filter predicates to produce populated results, and (3) injecting query-specific synthetic tuples in the sample, similar to the XData mutant-killing tool [34], to cover predicate boundary conditions – the complete details are in Section C.
+2.
+Logic-based Equivalence.
+Although verifying the equivalence of a general pair of SQL queries is NP-complete [7], a variety of logic-based tools (e.g.
+Cosette[12], SQL-Solver [17], VeriEQL [21], QED [40]) are available for proving equivalence over restricted classes of queries.
+Once the least-cost rewrite is identified, LITHE uses the recently proposed QED [40] and SQLSolver [17] since these two approaches were found to cover a larger set of queries compared to the alternatives.
+The advantage of such a logic-based approach is that it is definitive in outcome.
+Note that this rewrite has already passed the sampling-based tests described above.
 3.
 Result Equivalence on the Entire Database.
-Result equivalence could also be evaluated, in principle, on the entire database itself.
-Of course, this could turn out to be prohibitively expensive, especially if the queries themselves are time-consuming (e.g.
-due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions.
-Therefore, we leave this check as an optional choice for the DBA.
+If the logic-based test is inconclusive, result equivalence is evaluated on the entire database itself.
+The DBA may choose to prematurely terminate this test in case the checking time is found to be excessive.
+6.3 Regression Identification
+We use heuristics to predict whether the actual execution time of a given promising rewrite may turn out to be slower than the original query.
+The first heuristic is based on the SEER robust plan identification algorithm [19], which is predicated on the database engine providing a “force-plan” feature (e.g.
+[1]).
+Let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively.
+We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables.
+Then, by assigning appropriate values to these parameters, we construct queries that are located at the corners of the selectivity space.
+The plans $P^{U}$ and $P^{T}$ are forced at each of these corner locations, and if the rewrite’s cost is lower than the original at all of them, the rewrite is deemed to be robust.
+At this time, plan forcing is not available in all database engines.
+For such limited environments, we fall back to an alternative heuristic – we compare the query runtimes obtained on the series of executions carried out on sampled databases during semantic equivalence testing.
+If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust.
 7 Experimental Evaluation
@@
 As mentioned earlier, we focus on “slow queries” that take over 10 seconds to execute in their original form, operating in a cold-cache environment.
-LITHE has also been evaluated on other benchmarks, including DSB [13], ARCHER [44], JOB [19] and StackOverflow [27]. Due to space limitations, we defer their details to Section A.
+LITHE has also been evaluated on other benchmarks, including DSB [15], ARCHER [47], JOB [23] and StackOverflow [30].
+The overall performance characteristics were found to be similar to those presented here – due to space limitations, we defer their details to Section A.
 Testbed.
@@
 Metrics.
-For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed.
+For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with $>$ 1.5 speedup) could be constructed.
 Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e.
@@
 7.1.1 Estimated Cost
-LITHE produces a rewrite with a positive speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold.
+LITHE produces a rewrite with a positive cost speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold.
 Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5.
@@
 All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27.
-We were able to prove the rewriting correctness of 9 CPRs through the logic-based QED tool.
-The remaining passed our result equivalence tests on both the down-sampled and full databases.
-Furthermore, as a final check, we manually verified the correctness of the rewritten queries.
+Of these 27, we were able to formally verify 11 using the logic-based tools, whereas the remaining 16 passed our statistical tests.
+Furthermore, we also manually verified the correctness of these rewritten queries.
 A drill-down into the cost speedup performance at the granularity of individual queries is shown in Figure 8, which compares LITHE (orange bars) and SOTA (blue bars) on each of the 27 CPR queries – note that the cost speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale, and the queries are sequenced in decreasing order of LITHE speedup.
@@
 Figure 8: Execution Time Speedups
-Second, in most queries, the cost speedup of LITHE either exceeds or matches SOTA.
-In fact, there are several queries (e.g.
-Q45, Q25, Q4) where LITHE produces a highly beneficial CPR but SOTA simply returns the original query.
+Second, in most queries, LITHE’s cost speedup either exceeds or matches SOTA.
+In fact, for several queries (e.g.
+Q45, Q25, Q4), LITHE produces a highly beneficial CPR but SOTA returns the original query.
 Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query.
-And in Q88 and Q95, SOTA is marginally better than LITHE.
-At this stage, one could imagine that more rules could be added to LITHE to bring it on par with SOTA for queries like Q57.
-However, as discussed later in Section 8.3, such customized rules, while attractive at compile-time, run a significant risk of causing regressions at run-time.
-We have therefore selected only broad-brush rules for inclusion in LITHE.
-This conservative approach is justified below in the timing section, where we find that LITHE actually performed better than SOTA on these queries (Q57, Q88, Q95) wrt runtimes.
+And in Q88 and Q95, SOTA performs only marginally better.
+At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57.
+However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency.
+As the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) in terms of runtime.
 7.1.2 Execution Time
 Thus far, we had considered optimizer-estimated execution costs.
-We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA.
-We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 10000!
-Second, in almost all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
-There remain only a few instances (Q9, Q47) where SOTA is marginally better.
+We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA on a cold-cache environment.
+We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a large factor of 700!
+Second, in all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
 From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well.
-On the one hand, there is Q45 where the projected speedup of $10^{2}$ increases to a huge $10^{4}$ at runtime.
-On the other hand, the $10^{5}$ speedup for Q41 decreases to $10^{3}$.
+On the one hand, there is Q45 where the projected speedup of 40X increases to 700X at runtime.
+When run in a warm-cache environment, this speedup increases to a huge $10^{4}$ (see Section A).
+On the other hand, the $10^{5}$ speedup for Q41 decreases to a still reasonable 200X.
 But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100.
 However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites.
-Overall, LITHE produces more robust rewrites resulting in a high TSGM of 18.4, whereas SOTA provides a TSGM of 6.
+Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA provides a TSGM of 4.9.
 7.1.3 Reasoning
 As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries.
-For example, consider Q45 – the provided explanation correctly lists the main reasons for the huge $10^{4}$ speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc.
+For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700X speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc.
 Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@
 However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads.
-For instance, with Q11, the original query took nearly two hours to complete, whereas the LITHE rewrite executed in just 3 minutes.
+For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in just 3 minutes.
 Further, many applications tend to use a set of canned queries which are run thousands of times.
@@
 A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module.
-We explore techniques to improve their efficiency in the Section B.
+We explore techniques to improve their efficiency in Section B.
 7.3.2 Monetary Outlay
@@
 OptA OptB OptA OptB OptA OptB
-LITHE 7 8 3.6 4.8 1.5 2.4
-SOTA 3 6 2.1 3.5 1.3 1.7
+LITHE 12 9 3.6 4.1 2.1 1.9
+SOTA 3 5 1.5 1.3 1.4 1.2
 The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA.
-Despite the apparent lack of optimization headroom, LITHE still produces 7 and 8 CPRs resulting in a healthy CSGM of 3.6 and 4.8, respectively.
-Further, the TSGM provided by these rewrites are a useful 1.5 and 2.4, respectively.
-However, unlike the no-regression outcome seen with PostgreSQL, LITHE generates 3 CPRs on OptA, and 2 on OptB, which turn out to be regressions.
-So, although these commercial optimizers may have greater sophistication, they also suffer significant brittleness with regard to estimation accuracy.
-Interestingly, one regression on OptB was caused by the LLM bringing in, via a basic prompt, an unrestricted version of Rule 6, which is prone to generating poorly-estimated plans (see Section 8.3).
-Despite such occasional errors, LITHE provides good overall improvement in runtimes as indicated by the TSGM values.
-We also performed a preliminary study of LITHE on an internal real-world benchmark, whose data and query characteristics are more realistic than TPC-DS.
-Even on this hard benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with a CSGM of 2 and a TSGM of 3.3.
+Despite the apparent lack of optimization headroom, LITHE still produces 12 and 9 CPRs resulting in a healthy CSGM of 3.6 and 4.1, respectively.
+Further, the TSGM provided by these rewrites are a useful 2.1 and 1.9, respectively.
+Interestingly, although we did not observe any regressions with PostgreSQL, a few did surface in the commercial systems.
+Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites.
+As a case in point, a promising rewrite, as estimated by the optimizer, for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic, and the rewrite was abandoned.
+We also conducted a preliminary study of LITHE on an internal real-world benchmark with more realistic data and query characteristics than TPC-DS.
+Even on this benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with CSGM of 2 and TSGM of 3.3.
 The above results suggest LITHE has a useful role to play in industrial environments.
@@
 Table 8: Rewrite Performance on Masked Database.
-Approach # CPR CSGM
-TPC-DS Masked TPC-DS Masked
-SOTA 15 12 6.1 5.3
-LITHE 26 24 11.5 10.5
+Approach # CPR CSGM TSGM
+TPC-DS Masked TPC-DS Masked TPC-DS Masked
+LITHE 26 24 11.5 10.5 13.2 11.8
+SOTA 13 12 6.1 5.3 4.9 4.2
 7.5.2 LITHE on LLaMA
@@
 So, the message is that smaller models can also be fruitfully used in real-world environments.
-Table 9: LITHE Rewrite Performance with LLaMA.
+Table 9: LITHE Rewrite Performance with LLaMA
 # CPR CSGM TSGM
-LLaMA without MCTS 17 5.6 8.2
-LLaMA with MCTS 21 8.3 14.7
+LLaMA without MCTS 18 5.6 6.5
+LLaMA with MCTS 22 8.5 10.9
 8 The Road Ahead
@@
 Therefore, a fruitful area of future research could be achieving a middle-ground between the disparate world-views of LLMs and traditional optimizers.
-8.3 Revisiting Optimizer Plan Costing
-As highlighted in Section 7, there were instances of substantive differences between the promised speedup and that delivered at run-time.
-In fact, to the extent that speedups could even turn out to be regressions!
-This is due to the brittleness of optimizer plan costing in the new spaces explored by the LLM.
-As a case in point, during our LITHE design process, the prefiltering in Rule R6 (Table 1) had initially not been restricted to self-joins.
-It resulted in the number of CPRs (on PostgreSQL) being as high as 65, with an astonishing CSGM of 30.6!
-However, upon execution, most of these rewrites turned out to be regressions, which led to our inclusion of the restriction.
-But note that we are artificially modulating the rewrite, rather than fixing the plan costing module, which is the principled solution.
-In sum, while plan cost modeling has been a long-standing area of research, there is now even more reason given the new rewrite spaces to study this topic further – for instance, the operator cost model could be extended using calibration techniques similar to those advocated in [40], while the operator cardinality model could be improved with attention-based techniques [23].
 8.3 Agentic LLMs for Query Rewriting
@@
 8.4 Scope of Semantic Equivalence Tools
-As seen in the experiments section, the coverage of logic-based query equivalence testing on industrial-strength queries is rather limited.
+As seen in the experiments section, logic-based query equivalence testing covers industrial-strength queries only to a limited extent.
 On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness.
@@
 Most of the recent work on SQL query rewriting is rule-based [48, 42, 9, 43, 11, 31].
-For instance, WeTune [38] employs a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules.
+For instance, WeTune [42] uses a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules.
 While this approach can generate a large set of new rewrite rules, it often fails in coming up with transformation rules for complex queries due to the computational overheads of verifying rule correctness.
 As such, it is unable to rewrite any of the TPC-DS queries [18, 28].
-Learned Rewrite [45] uses a set of existing rewrite rules from Calcite [9] and aims to learn the optimal subset of rules along with the order in which they must be applied.
-Given that the search space for possible rewrites is exponential in the number of applicable rules, an MCTS scheme is used to efficiently navigate this space and find the rewritten query with maximum cost reduction.
+Learned Rewrite [48] uses existing Calcite [10] rules and aims to learn the optimal subset of rules along with the order in which they must be applied.
+Since the rewrite search space grows exponentially with the number of rules, it uses MCTS scheme to efficiently navigate this space and find the rewritten query with maximum cost reduction.
+This is in contrast to LITHE where MCTS is used to search over the output token space of an LLM in order to prevent LLM hallucinations.
+Specifically, our use of MCTS ensures that the LLM correctly follows the prompt instructions.
 LLM-R2 [27] is also rule-based but takes a different approach to identify the order for rewrite rule applications: it uses an LLM to find the best Calcite rules and the order in which to apply them to improve the query performance.
@@
 LLM-based rewriting.
-GenRewrite [25] is the first LLM-based approach that tries to use the LLM itself for end-to-end query rewriting.
+GenRewrite [28] is the first LLM-based approach to use the LLM for end-to-end query rewriting.
 Instead of using predefined rules from Calcite [10], they employ the LLM to create Natural Language Rewrite Rules (NLR2s) to be used as hints, and perform several iterations of prompting to get the rewritten query.
@@
 Appendix
-Appendix A Rewrite Quality (Other Benchmarks)
+Appendix A Additional Experiments
+A.1 Rewrite Quality (TPC-DS) Contd.
+Figure 10 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA for the queries run on a warm-cache environment.
+Figure 10: Execution Time Speedups in a warm-cache scenario.
+As expected, the speedups often significantly increase when compared to the cold-cache scenario.
+For instance, the speedup for Q45 increases from 700X to over $10^{4}$.
+Overall, the TSGM of LITHE increases to 17.1 (from 13.2), while that of SOTA increases to 5.7 (from 4.9).
+A.2 Rewrite Quality (Other Benchmarks)
 We also evaluated LITHE on the following set of benchmarks.
@@
 The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent.
-However, there is a price to be paid – the CPR is reduced to 24 and the CSGM comes down to 8.5.
+However, there is a price to be paid – the CPR is reduced to 23 and the CSGM comes down to 8.5.
 In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
@@
 Appendix C Query Equivalence Testing via Sampling
-If the original query is not within the scope of the logic-based QED tool, we alternatively use a sampling-based approach to test equivalence.
+We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline.
 The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results.
@@
   2. 2.
-    To differentiate between inner and outer joins, we insert into the sample database, rows with appropriate values on the relevant columns such that they produce NULL values for outer joins.
-  3. 3.
     Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result.
@@
     Note that these modifications only change the selectivity of the query, but not its semantics.
+  3. 3.
+    To cover predicate boundary conditions, we inject query-specific synthetic tuples into the sample—similar to the approach used by the XData mutant-killing tool [34]—in order to reduce the likelihood of non-equivalent queries producing the same results.
 Appendix D Detailed MCTS Algorithm
@@
 10:    while $len(n_{cur}.children)>0$ do
-11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$
+11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
 12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$

### v4

@@
 We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms.
-The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer.
-For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
+The results show significant performance improvements for slow queries, over both SOTA rewriters and the native optimizer.
+For instance, with TPC-DS on PostgreSQL, the GM of runtime speedups was a high 13.2 over the native optimizer, whereas SOTA only gave 4.9.
 Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise query performance.
 1 Introduction
-The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools.
+The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies [51, 2, 6], especially when machine-generated by modeling software such as ORM tools (e.g., Entity Framework [37], Hibernate [27]).
 Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans.
 However, in practice, they are often led astray by complex representations, resulting in poor response times.
-SELECT t.Key, sum(t.Rating) AS PostRating,
-(SELECT sum(b0.Rating)
-FROM (SELECT p0.PostId, p0.BlogId, p0.Content,
-p0.CreatedDate, p0.Rating, p0.Title,
-b1.BlogId AS BlogId0, b1.Rating AS Rating0,
-b1.Url, p0.day AS Key
-FROM Posts AS p0 INNER JOIN Blogs AS b1
-ON p0.BlogId = b1.BlogId
-WHERE b1.Rating > 5) AS t0
-INNER JOIN Blogs AS b0 ON t0.BlogId = b0.BlogId
-WHERE t.Key = t0.Key ) AS BlogRating
-FROM (SELECT p.Rating, p.day AS Key
-FROM Posts AS p INNER JOIN Blogs AS b
-ON p.BlogId = b.BlogId
-WHERE b.Rating > 5) AS t
-GROUP BY t.Key;
 Figure 1: Complex SQL Representation
-SELECT p.day AS Key, SUM(p.Rating) AS PostRating,
-SUM(b.Rating) AS BlogRating
-FROM Posts AS p INNER JOIN Blogs AS b
-ON p.BlogId = b.BlogId
-WHERE b.Rating > 5
-GROUP BY p.day;
 Figure 2: Lean Equivalent Query
-As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs.
+As a case in point, consider the blog-processing query [6] shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs.
 When this seemingly complex query is given to the current PostgreSQL engine (v16) [3], its execution plan essentially mimics the hierarchical structure of the query.
@@
 However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
-For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives.
-For this context, a viable S2S query transformer should satisfy the following criteria: (1) A wide range of queries should be covered in its scope; (2) The transformed query should be semantically equivalent to the original; (3) The rewrite should ideally improve performance, but at least not cause regression; and (4) The transformation overheads must be practical for deployment.
+For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools for recommending performant alternatives.
+A viable S2S query transformer should cover a wide range of queries and ensure the rewrite is: (1) semantically equivalent to the original; (2) ideally performance-beneficial, but at least not causing regression; and (3) incurring overheads that are practical for deployment.
 Prior Work
@@
 Gen-Rewrite [34]) approaches have been proposed for S2S transformations.
-A recent vision paper [49] also advocates the use of LLMs for such query rewriting.
-The state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting.
-However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e transforming the query structure itself).
-In this paper, we address these lacuna and amplify the effectiveness of query rewriting.
-In particular, our query space-based transformations provide greater scope for candidate rewrites since only intent, and not implementation, is expressed, and this multiplicity of rewrites in turn results in a broader coverage of the underlying plan space.
-Figure 3: High-level architecture of LITHE
+These state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting.
+However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e.
+transforming the query structure itself).
+In this paper, we address these lacuna and substantively amplify the effectiveness of query rewriting.
 The LITHE Rewriter
-We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention.
-As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance.
+We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries.
+As illustrated in the architectural diagram of Figure 3, LITHE takes the user query $Q^{U}$ as input and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement of $Q^{T}$, in terms of optimizer estimated cost; (b) a verification label indicating the semantic equivalence mechanism (provable or statistical) between $Q^{T}$ and $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance.
 Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$.
 Note that having the DBA in the loop is a common practice in commercial query advisory systems [17].
-Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
-Basic Prompts.
-We begin with a suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs.
-Interestingly, we find that more information is not necessarily better wrt rewriting quality, and that the best prompt granularity is query-specific.
-Moreover, this basic prompt ensemble was found to itself deliver performance similar to the SOTA techniques.
-Database-sensitive Prompts.
-To help the model adapt to different query patterns and structures, we next introduce rules in the prompts.
+To help the LLM adapt to different query patterns and structures, we introduce database-aware rules into the prompts, which make use of the rich metadata (e.g., logical schema, selectivities) available in database environments (Section 4).
 Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries.
-This is in contrast to the hardwired and narrow rule application mechanisms (e.g.
-Calcite [10] rules) typically used in existing rewrite systems, which operate in plan space.
-In particular, we work with two classes of rules – 1.
-Redundancy Removal Rules: These rules eliminate repeated and redundant computations of the same output; and 2.
-Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts.
-To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context.
-As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
-Token Probability Driven Rewrites.
-The above rewrite options are restricted to the standard prompt interface.
-We also leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence.
+As shown later in our experiments, this is a powerful mechanism for ensuring performant rewrites across database environments.
+Furthermore, we leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence.
 Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process.
-To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation.
-MCTS has previously been used to improve ordering of query rewrite rules [48] in the plan space, but our usage is new – it is geared to guide the LLM in exploring rewrites in the query space.
-Semantic Equivalence Labels.
-After generating a rewrite, it is imperative to check for semantic equivalence to the original query.
-Unfortunately, the general problem of proving query equivalence is NP-complete [7].
-We therefore take a twofold approach to check for query equivalence.
-First, the queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database.
-While these checks do not suffer from false negatives, they may (rarely) incur false positives.
-These tests can thus quickly weed out obviously non-equivalent rewrites.
-A final check is then performed using logic-based equivalence-proving tools that provably work for a restricted classes of queries.
-Specifically, LITHE uses the recently developed QED tool [40] and SQLSolver [17] to verify rewrite correctness for queries within their scope.
-Whereas, for out-of-scope queries, result equivalence on the original database itself is used to test the equivalence.
-After carrying out the above exercise, LITHE provides a verification label of semantic equivalence – provable if verified by a logic-based tool, or statistical otherwise.
-Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
-Regression Identification.
-Finally, it is well-known that in practice there are often material discrepancies between optimizer cost predictions and actual execution times [29].
-So it is possible that a promising rewrite as per the optimizer may actually regress at run-time as compared to the original formulation.
-To help guard against such possibilities, we have implemented heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
-Experimental Evaluation
-Our first set of experiments to evaluate LITHE’s performance is carried out on the industry standard TPC-DS benchmark [13], hosted on the PostgreSQL platform with GPT-4o used as the LLM.
-The evaluation focuses on slow queries taking more than a threshold time to complete.
-We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [48], LLM-R2 [27], GenRewrite [28], as well as a baseline LLM prompt [28]).
-The primary metrics are (a) reductions in optimizer-estimated costs, (b) run-time speedups, and (c) rewriting overheads.
-For LLM-based techniques, the number of tokens used is also monitored since the financial charges for LLM usage are typically dependent on this number.
-To understand the independent utility of the various components of LITHE, a systematic ablation study is carried out.
-Generalizability.
-In our second stage of experiments, we evaluate generalizability of the above outcomes in a variety of new scenarios, including (a) commercial database engines, (b) unseen database schemas, and (c) alternative LLM platforms.
+To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation (Section 5).
+LITHE also incorporates a suite of logic-based and result equivalence-based semantic equivalence tests to verify the correctness of the rewrites (Section 6.2).
+Finally, given that there are often material discrepancies between optimizer cost predictions and actual execution times [35], LITHE includes heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
+Figure 3: High-level architecture of LITHE
 Results.
-Our experiments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract costs.
-In particular, for TPC-DS, LITHE constructed “highly productive” ($>1.5x$ estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number.
+Our experiments on benchmark environments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract (optimizer-estimated) costs.
+In particular, for TPC-DS, LITHE constructed “highly productive” (>1.5x>1.5x estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number.
 Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
-We also evaluated whether the above cost reductions translated into real execution speedups.
+We also evaluated whether the projected cost reductions translated into real execution speedups.
 Here, we find that LITHE is indeed often substantively faster at run-time as well.
@@
 Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
-Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
 Contributions
@@
   5. 5.
-    Evaluates rewrite quality over a broad range of database environments, demonstrating substantial benefits over both SOTA and the native optimizer.
+    Evaluates rewrite quality over a range of database environments, demonstrating substantial performance benefits.
   6. 6.
@@
 2 LITHE Overview
-The LITHE architecture, illustrated in Figure 3, consists of the following five-stage pipeline:
+The LITHE architecture (Figure 3) comprises a five-stage pipeline:
 1.
@@
 The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite.
-The LLM output’s syntax is then verified with the database engine’s parser– if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively (within a threshold) until a valid SQL query is obtained.
+The LLM output’s syntax is then verified with the database engine’s parser – if invalid, the error is returned to the LLM via an updated prompt asking for a correction.
+This goes on iteratively (within a threshold) until a valid SQL query is obtained.
 This process is repeated for each of the different prompts described in Sections 3 and 4, and the resulting syntactically-valid rewrites are fed to the query costing component.
@@
 For this context, we define a Cost Productive Rewrite (CPR) as a rewrite that improves a slow query’s performance by at least 1.5 times wrt the optimizer-estimated cost – this aggressive choice of threshold is so that: (a) there is enough headroom in the optimizer prediction that a runtime regression is unlikely; and (b) the benefits of the rewrite substantively outweigh the transformation overheads.
-The overall benefit provided by a rewriting tool is quantified by the number of CPR obtained on the slow query workload.
-Additionally, we also measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites.
+The overall benefit of a rewriting tool is quantified by the number of CPR obtained on the slow query workload.
+Additionally, we measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites.
 Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites.
+These speedups are measured relative to the original query on the native optimizer.
 2.2 Query Micro-benchmark
@@
 Later, in Section 7, we extend the evaluation to complete benchmarks.
-Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sec 345). The TSGM performance is subsequently discussed in Section 7.
+Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sections 345). The TSGM performance is subsequently discussed in Section 7.
 3 Basic Prompts
@@
 Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once.
-The idea is to break down the complex instructions given in Prompt 3 into digestible steps that help the LLM focus on individual tasks.
+This breaks down the complex instructions of Prompt 3 into digestible steps that help the LLM focus on individual tasks.
 Performance
@@
 As discussed above, basic prompting needs to be improved on two fronts: (1) Ensuring productive rewrites where feasible; and (2) Maximizing the impact of these productive rewrites.
-To address these issues, we incorporate database domain knowledge.
+Our discussions with industry experts revealed that query inefficiencies are most commonly attributed to redundant operations within the queries themselves.
+In fact, contemporary (non-LLM) optimizers often try to reduce redundancies during the initial logical optimization phase, prior to physical plan enumeration (e.g., redundant join elimination [9]).
+Furthermore, our analysis of execution plans for slow queries revealed that the optimizer often fails to substitute filter-related operators, even when beneficial to do so.
+To address these issues, we incorporate database domain knowledge into the prompts.
 Specifically, we design a one shot-based prompting template, augmented with a set of database-aware rewrite rules.
-The rules are based on common practices followed by DBAs that are widely applicable, and augmented with precise instructions and useful examples to help guide the LLM in the rewriting process.
 As a proof of concept, we explore two categories of rewrites here: (a) Rules that eliminate redundancy in the input queries; and (b) Predicate selectivity-based rules that implicitly guide, via query space reformulations, the query optimizer towards efficient query execution plans.
 Of course, this basic set of rules can be expanded further, but as shown by our experiments, even this minimal set is capable of delivering substantive improvements over a broad set of database environments.
+The templates used for these prompts (Figure 5) apply only a single rule.
+This was a conscious design choice because LLMs can be overwhelmed by excessive information given in monolithic form [46].
+A related question is whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied, thereby benefiting queries with multiple types of redundancies.
+However, the combinatorially large number of rule permutation sequences makes the selection process appear infeasibly expensive.
+Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best improvement.
+Despite this restrictive policy, highly performant rewrites are obtained in our evaluation (Section 7).
 4.1 Redundancy Removal
@@
 Figure 5: Templates for Database-sensitive Prompts
-The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are available in the Section E.
-Note that this prompt template allows for only a single rule to be present in the prompt.
-This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form.
-Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
+The template for such rule-based prompts is shown in Figure 5(a) and includes an example rewrite to demonstrate the rule application to the LLM.
+Our empirical observation is that the precise wording of the rule instructions is not significant.
+Instead, what matters are the example rewrites, which ensure the LLM applies the rules effectively, and does not go off on tangential trajectories.
+The specific examples used with our rules are listed in Section D.
 Performance
@@
 Prompt # CPR CSGM
-[…]
-A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on.
-This approach would benefit queries with multiple types of redundancies.
-However, it also introduces significant computational overheads due to the vast number of possible rule application orders, making it expensive to explore.
-So, for simplicity, we have chosen to process them individually rather than cumulatively.
+Basic Prompts $\bigcup$ {R1, $\ldots$, R4} 7 6.85
 Figure 6: Example Queries illustrating Rule 5
 4.2 Selectivity-based Guidance
-We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects.
-For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values.
-Rule R5 basically encodes this argument as a rule in Table 2.
-Similarly rule R6, which pre-filters tables that are involved in self-joins and have low selectivity filter and/or join predicates, is also used to guide a rewrite.
-Note that a specific instruction to not create explicit joins had to be added in rule R6.
+We now consider rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects.
+For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively.
+Here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values.
+This observation is encoded in Rule R5 of Table 2.
+Similarly rule R6, which pre-filters tables that are involved in self-joins and with low selectivity filter and/or join predicates.
+Note that a specific instruction to not create explicit joins is added in rule R6.
 This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
@@
 Prompt # CPR CSGM
-[…]
+Basic Prompts $\bigcup$ {R1, $\ldots$, R6} 9 10.57
 We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3.
@@
 5 Token Probability Driven Rewrite
-A key challenge with LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up.
-This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs [20].
-In order to have a robust approach for such cases, we take inspiration from the code generation literature [46].
-Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
-As mentioned earlier, MCTS has been previously used for ordering rule applications in query rewrites [48], whereas our goal is to guide the LLM in exploring query rewrites.
+A key challenge with LLMs is “hallucinations” – responses that range from being mildly incorrect to completely made up.
+This is often due to the output tokens having low confidence, which “confuse” the LLM and generate suboptimal outputs [23].
+To have a robust approach for such cases, we take inspiration from the code generation literature [55].
+Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite and performance improvements.
+While MCTS has been previously used for ordering rule applications in query rewrites [57], our goal is to guide the LLM in exploring these rewrites.
 This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [41].
 The root node of the tree corresponds to the initial prompt.
+Figure 7 illustrates one such example tree.
 An edge from a parent node to a child represents a possible token generated by the LLM and is associated with a value denoting the probability of generating this token given the path taken thus far.
+This is illustrated by the text (token) and number (probability), on the left and right side respectively, of an edge in Figure 7.
 Here, each edge can be considered as an action of the MDP.
-A node is considered terminal if the incoming edge corresponds to the “;” token, signalling end of the textual query.
+A node is considered terminal if the incoming edge corresponds to the “;” token, signaling end of the textual query.
 The state of a node $n$ is represented by the partial rewrite created by following the path from the root to $n$ – it is obtained by concatenating the tokens on this path.
 The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query.
+For example, the state of Node 4 in the example decision tree shown in Figure 7 is the partial rewrite ‘‘select * from’’.
 Given the hundreds of thousands of tokens in LLM vocabularies, it may be very expensive (both financially and computationally) to construct the entire tree.
@@
 Algorithm 1 Token-augmented Rewrite
-1:$root$          # Start State
+1:rootroot    # Start State
 2:$k$      # Maximum number of child node expansions
-3:$\theta$              # Probability threshold for node expansion
-4:$iter_{max}$     # Maximum number of iterations
-5:$Potential$, $visits$, $V$leftarrow empty Map
+[…]
+[…]
+5:PotentialPotential, visitsvisits, $V$ $\leftarrow$ empty Map
 6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
-7:    $visits[root]$leftarrow $visits[root]+1$
-8:    $n_{cur}$leftarrow $root$
+7:  $visits[root]$ $\leftarrow$ $visits[root]+1$
+8:  $n_{cur}$ $\leftarrow$ rootroot
 9:  # Stage 1: Selection
 10:  while $len(n_{cur}.children)>0$ do
-11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
-12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
+[…]
+12:   $visits[n_{cur}]$ $\leftarrow$ $visits[n_{cur}]+1$
 13:  # Stage 2: Expansion
 14:  $expand\leftarrow$ True
-15:    while $expand$ and ‘;’ $\notin n_{cur}.state$ do
-16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
+15:  while expandexpand and ‘;’ $\notin n_{cur}.state$ do
+16:   $tokens_{next}$, $P_{next}$ $\leftarrow$ Model($n_{cur},k$)
 17:   if $P_{next}[0]\leq\theta$ then
 18:     for $token\in tokens_{next}$ do
-19:               $n_{new}\leftarrow$ new $Node$ with State $n_{cur}.state\cdot token$
+19:      $n_{new}\leftarrow$ new NodeNode with State $n_{cur}.state\cdot token$
 20:      Append $n_{new}$ to $n_{cur}.children$
@@
 25:  $query\leftarrow$ GreedyExpand($n_{cur}$)
-26:    $v\leftarrow$ ComputePotential($query$)
+26:  $v\leftarrow$ ComputePotential(queryquery)
 27:  $Potential[query]\leftarrow v$
@@
 5.1 MCTS Search Process
-The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
+The search procedure is codified in Algorithm 1, and comprises four stages – Selection, Expansion, Simulation and Back Propagation – described below, which are repeated for $iter_{max}$ iterations.
 1.
-Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure).
+Selection: The first stage identifies the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure).
 Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1).
@@
 The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often.
-Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process.
+Here, $P_{LLM}$ represents the next token probability, which can be queried from the LLM by enabling the logprobs option in the API, and $visits[n]$ is the number of times $n$ has been visited during the search process.
 $\beta$ is a function that controls the balance between exploration and exploitation.
@@
 $\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
+Figure 7: Sample decision tree traversal (Algorithm 1)
+For example, consider the state of the tree in Figure 7 when Nodes 5 and 15 are the current unexpanded nodes.
+At this juncture, the selection procedure will use the UCB values of these two nodes to choose which node to expand next.
 2.
@@
 To make the expansion tractable, multiple child nodes are added only if the probability of the highest token falls below a threshold $\theta$ (Line 13).
-In other words, when the highest token probability is below $\theta$, it means that the LLM itself is unsure of what the next token should be and therefore it is worth exploring additional options.
-On the other hand, if the highest token probability is greater than $\theta$, then the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e., completes a query rewrite).
+Otherwise, the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e.
+completes a query rewrite).
+In our implementation, $k$ and $\theta$ are set to 2 and 0.7, respectively.
+Since branching occurs only when this relatively high $\theta$ threshold is breached, most nodes have only one child, resulting in long linear sequences.
+This design, combined with at most two nodes being processed during a branch, significantly reduces the effective fan-out of the tree.
+Moreover, it ensures that MCTS focuses computational resources on the brittle segments of the rewrite, and does not waste them on robust modifications.
+For example, when the root node in Figure 7 is first processed, then Nodes 1, 2, 3 and 4 are expanded and created sequentially since the token probabilities are higher than $\theta$ (Lines 24–26 of the while loop).
+Only when Node 4 is reached, two new nodes (5 and 15) are created as its children.
 3.
-Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$.
-This node is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21).
-Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23).
-For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides w.r.t the original query.
+Simulation: The node $n_{cur}$ is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21).
+The rewritten query obtained from this greedy expansion is then used to compute the potential $V(n_{cur})$ (Line 22) as follows: for a valid rewrite, $V(n_{cur})$ represents the speedup it provides w.r.t.
+the original query.
 However, if invalid (i.e.
 syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value.
-After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
+After every simulation, the rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$ (Line 23).
+For instance, to compute $V(n_{15})$ in Figure 7, the blue path in the tree is greedily expanded to identify its rewrite potential.
+That is, the partial rewrite “select * from B left outer join A;” is used to evaluate $V(n_{15})$.
 4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is back-propagated to all its ancestor nodes, and their values are updated iff the new value is higher than the existing value.
@@
 While these gains over just prompting may seem marginal, we wish make two observations: (1) MCTS extracts maximum improvement whererever possible, and (2) As shown in Section 7.5.2, its impact becomes significant for smaller LLMs like LLaMA.
+Discussion.
+Our approach implicitly assumes that LLMs do possess some degree of knowledge about query optimization – for example, when presented with a poorly written query versus an optimized one, the LLM is usually capable of identifying why the optimized version would execute better.
+This expertise arises from the extensive training corpus of LLMs, which includes a large repository of books, papers, and articles on query optimization and execution.
+However, this core knowledge is not fully sufficient to provide good rewrites, and there is a clear need for augmentation with explicit query-specific guidance.
+This is particularly so when there are multiple plausible optimization paths, or when the LLM may be subtly misled – this is the reason we leverage MCTS to systematically explore alternatives based on token probabilities.
+Further, there is explicit evaluation via the database optimizer of the expected improvement in quantitative performance.
+This separation ensures that our method leverages the LLM’s generative capabilities to posit rewrites while relying on an external ground-truth signal (i.e.
+the database optimizer) to confirm rewrite quality.
 6 Implementation Choices
@@
 These settings were determined after an empirical evaluation of the various trade-offs, providing a robust balance between efficiency and quality.
-Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2).
+Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2) – if this threshold is crossed, we drop the rewrite.
 6.2 Query Equivalence Testing
 We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite.
-Note that, as mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems[14].
+As mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems [17].
 1.
@@
 Cosette[15], SQL-Solver [20], VeriEQL [26], QED [49]) are available for proving equivalence over restricted classes of queries.
-Once the least-cost rewrite is identified, LITHE uses the recently proposed QED [40] and SQLSolver [17] since these two approaches were found to cover a larger set of queries compared to the alternatives.
+Once the least-cost rewrite is identified, LITHE uses QED [49] and SQLSolver [20] since these two approaches covered a larger set of queries compared to the alternatives.
 The advantage of such a logic-based approach is that it is definitive in outcome.
@@
 [1]).
-Let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively.
+This approach is based on the notion of replacement safety, wherein the replacement plan is always either better or within a small sub-optimality factor of the original optimizer-recommended choice.
+It was theoretically shown that a safe plan replacement at the perimeter of the selectivity space is sufficient to infer safe replacement in the interior of the space as well.
+The proof used a generic characterization of plan cost functions and the first and second derivatives of these functions at the perimeter.
+As a matter of efficiency, it was empirically found that instead of the entire perimeter, it was usually sufficient to consider safety just at the corners.
+We use the above result to design a regression test as follows: let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively.
 We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables.
@@
 If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust.
+We found this extrapolation heuristic to be effective in our experiments due to two reasons: First, the sampled databases, while relatively small compared to the original, are still large in absolute terms (1 GB or more).
+Second, since most rewrites are related to redundancy removals, a relative performance improvement obtained due to such a rule holds true irrespective of the data sizes.
 7 Experimental Evaluation
@@
 LITHE has also been evaluated on other benchmarks, including DSB [18], ARCHER [56], JOB [29] and StackOverflow [36].
-The overall performance characteristics were found to be similar to those presented here – due to space limitations, we defer their details to Section A.
+The overall performance characteristics were found to be similar to those presented here – see Section A for details.
 Testbed.
-The majority of our experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA.
-Variations on this platform are considered in Sections 7.4 and  7.5.
+The experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660, 32 GB RAM, and 12 TB HDD, running Ubuntu 22.04 LTS.
+A majority of the experiments used PostgreSQL v16 database engine and GPT-4o LLM for both LITHE and SOTA.
+Commercial engines and other LLMs are considered in Sections 7.4 and  7.5.
 Metrics.
-For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with $>$ 1.5 speedup) could be constructed.
+For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed.
 Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e.
 CPRs arising from either LITHE or SOTA).
+Cost speedups are computed relative to the native optimizer, and runtime speedups are measured as the ratio of original query runtime to rewritten query runtime in a cold-cache environment.
 From the investment perspective, we measured the average rewrite time per query, and additionally for the LLM-based techniques, the number of tokens used in the rewrite process.
@@
 7.1.1 Estimated Cost
-LITHE produces a rewrite with a positive cost speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold.
+LITHE produces a rewrite with a positive cost speedup (¿ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold.
 Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5.
@@
 The vertical dotted line at $1$ represents the normalized baseline cost of the original query with the native optimizer, while the vertical line at $1.5$ is the CPR threshold.
-We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE.
-This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
+We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups over the native engine – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE.
+This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter efficiently checks result existence in an inner subquery since it removes the costly aggregation function.
 Figure 8: Plan Cost Speedups
 Figure 9: Execution Time Speedups
-Second, in most queries, LITHE’s cost speedup either exceeds or matches SOTA.
+Second, in as many as 15 queries, LITHE’s cost speedup substantively exceeds SOTA, while in the remainder it largely matches SOTA.
 In fact, for several queries (e.g.
-Q45, Q25, Q4), LITHE produces a highly beneficial CPR but SOTA returns the original query.
-Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query.
-And in Q88 and Q95, SOTA performs only marginally better.
+Q45, Q25, Q4) LITHE produces a highly beneficial CPR but SOTA returns the original query.
+The only case where SOTA is appreciably better is Q57 where it projects a large speedup but LITHE settles for the original query.
+And in Q88 and Q95, SOTA performs marginally better.
 At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57.
 However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency.
-As the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) in terms of runtime.
+Moreover, as the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) at runtime!
+We also analyzed queries where LITHE was unable to provide CPRs to determine whether there were common patterns that led to unsuccessful outcomes.
+Our audit indicated that the primary reasons were: (a) Structural Simplicity – for instance, flat SPJ queries without nesting that are already conducive to current optimizers (e.g.
+Q3, Q15, Q52), or (b) Structural Tightness – for instance, queries without repeated computations where the scope for improvement is inherently limited (e.g.
+Q7, Q19, Q22).
+Further, some failures may be due to the specific LLM being used, and not the queries themselves (see Section 7.5.2).
+Therefore, one could consider implementing an ensemble of diverse LLMs to gain enhanced beneficial coverage.
 7.1.2 Execution Time
-Thus far, we had considered optimizer-estimated execution costs.
-We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA on a cold-cache environment.
-We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a large factor of 700!
-Second, in all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
+Thus far, we had considered optimizer-estimated costs.
+We now move to wall-clock runtimes for query executions – Figure 9 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA.
+We again observe that there are indeed several queries where substantial time benefits are achieved by the rewrites over the native engine, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a huge factor of 700!
+Further, for 12 of the 26 queries, LITHE significantly outperforms SOTA – as a case in point, by 14 times in Q11.
+In the remainder, it matches SOTA, including as mentioned above, queries where SOTA’s optimizer costs were better.
 From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well.
-On the one hand, there is Q45 where the projected speedup of 40X increases to 700X at runtime.
-When run in a warm-cache environment, this speedup increases to a huge $10^{4}$ (see Section A).
-On the other hand, the $10^{5}$ speedup for Q41 decreases to a still reasonable 200X.
+On the one hand, there is Q45 where the projected speedup of $40$ increases to $700$ at runtime, whereas on the other, the $10^{5}$ speedup for Q41 decreases to $200$.
 But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100.
-However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites.
-Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA provides a TSGM of 4.9.
+However, the good news is that with LITHE, although the runtime speedups did not always match the projections, regressions were not encountered among the CPR rewrites thanks to the sampling-based checks.
+Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA only delivers TSGM of 4.9.
 7.1.3 Reasoning
 As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries.
-For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700X speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc.
+For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700700x speedup as “the use of CTEs”, “Reduction in data volume early on” due to pre-filtering based on join predicates, etc.
 Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
@@
 When the statistics-infused Rules (R5, R6) are added, this number jumps to 25.
-Finally, MCTS produces an additional CPR and also improves the cost speedup of a pre-existing CPR.
+Finally, MCTS in isolation (with Prompt 1 as the seed prompt) produces only 11 CPRs.
+However, by using the best prompt as seed, an additional CPR is obtained and the cost speedup of a pre-existing CPR is also improved.
 7.2.2 Database-Sensitive Prompts.
 Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance.
-The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads.
+The motivation is that it would relieve us from using the database-sensitive rules R1–R6 which incur significant computational and financial overheads.
 When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules.
@@
 7.3.1 Transformation Time
-The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes.
-However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads.
-For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in just 3 minutes.
+We first look at the end-to-end time required to perform a rewrite (i.e.
+to run the entire pipeline shown in Figure 3).
+Table 6 shows the average processing time per CPR query, where we see that the LITHE rewrite process takes a few minutes.
+However, note that this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads.
+For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in under 3 minutes.
 Further, many applications tend to use a set of canned queries which are run thousands of times.
@@
 Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites.
-A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module.
+A drill-down showed the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module.
 We explore techniques to improve their efficiency in Section B.
@@
 Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites.
-As a case in point, a promising rewrite, as estimated by the optimizer, for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic, and the rewrite was abandoned.
-We also conducted a preliminary study of LITHE on an internal real-world benchmark with more realistic data and query characteristics than TPC-DS.
-Even on this benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with CSGM of 2 and TSGM of 3.3.
+As a case in point, a promising rewrite estimated by one of the commercial optimizers for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic (Section 6.3), and the rewrite was abandoned.
 The above results suggest LITHE has a useful role to play in industrial environments.
 From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
-7.5 Dependence on LLM (Training/Model)
-7.5.1 Masked Database
+7.5 Dependence on LLM
+7.5.1 Impact of Training
 An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain.
-To investigate this issue, we created a masked version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their contents.
-We then constructed rewrites for the CPR queries (after syntactic changes to reflect the new masked schema) on this version.
-The results are shown in Table 8, and we observe that the performance profiles only marginally decrease for both LITHE and SOTA, testifying to their robust applicability.
-Table 8: Rewrite Performance on Masked Database.
-Approach # CPR CSGM TSGM
-TPC-DS Masked TPC-DS Masked TPC-DS Masked
-LITHE 26 24 11.5 10.5 13.2 11.8
-SOTA 13 12 6.1 5.3 4.9 4.2
-7.5.2 LITHE on LLaMA
+To investigate this issue, we ran LITHE on two datasets that were confirmed to be unknown to the GPT-4o version used in our evaluation.
+The first is the recently released Football benchmark [24] for Text-to-SQL evaluation.
+From the query suite, which largely consists of simple SPJ and Union queries, we selectively picked the few complex queries.
+Further, we created our own complex queries to include in the test suite.
+On this enriched workload, LITHE on PostgreSQL produced 11 CPRs with CSGM of 12 and TSGM of 1.5, while SOTA gave 6 CPRs with CSGM of 3 and TSGM of 1.1.
+The second dataset is a proprietary customer benchmark used extensively by an industrial development team.
+For this environment, LITHE on a commercial engine produced 8 CPRs with a CSGM of 2 and a TSGM of 3.3.
+Note that these improvements are achieved despite this engine having been fine-tuned on this workload over an extended period.
+Moreover, the numbers are consistent with those reported for TPC-DS in Section 7.4.
+7.5.2 Impact of Model
 In our concluding experiment, we evaluated LITHE’s performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which may have several hundred billion parameters [8].
@@
 To make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
-For this environment, Table 9 shows LITHE’s performance on the TPC-DS benchmark with and without MCTS.
+For this environment, Table 8 shows LITHE’s performance on the TPC-DS queries with and without MCTS.
 Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support.
@@
 LLaMA with MCTS 22 8.5 10.9
+Finally, we also evaluated LITHE on Gemini 2.5 Flash.
+Here, LITHE produced 17 CPR with CSGM of 7.2 and TSGM of 5.8.
+Although not as strong as GPT-4o, these results highlight the potency of LITHE’s recommendations across a range of LLM platforms.
 8 The Road Ahead
@@
 These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines.
+Specifically, rewrites in the query space appear to guide the optimizer to explore fresh regions of the plan space not evaluated with the original query formulation.
 Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [11].
+We plan to explore these space relationships and options in our future work.
 8.2 Rewrite Migration to Optimizer
@@
 Therefore, it would be a useful exercise to try and distill fresh optimization rules from these instances, leveraging the extensibility features of contemporary optimizers [19] to facilitate their incorporation in existing systems.
-Figure 9: Rewritten TPC-DS Q90
 On the flip side, there appears to be an “impedance mismatch” against such integration for certain classes of rewrites.
@@
 The rewrite, however, extracted all relevant rows in one shot and computed the ratio using CASE statements – encoding such transformations as generic rules in the optimizer appears challenging, given the combinatorial ways in which such transformations can occur.
+Figure 10: Rewritten TPC-DS Q90
 Therefore, a fruitful area of future research could be achieving a middle-ground between the disparate world-views of LLMs and traditional optimizers.
@@
 An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment.
-This would allow a “LITHE agent” to leverage a memory store to log one or more of: (1) query execution times; (2) memory usages of slow queries; (3) prior interactions with users; and (4) useful/new rewrite rules learnt over time.
-Additionally, the agent could be given database engine access so that it can directly request metadata such as query execution plans, database statistics, live query analyzers, etc.
-By following a structured workflow – planning, acting, observing, and refining – the agent would then iteratively enhance query performance.
-This adaptive workflow supports continuous learning, so that query rewriting can become more efficient over time.
+This would allow a “LITHE agent” to leverage database-related “tools” such as the query optimizer’s cardinality estimation and cost estimation modules, as well as rewrite-related tools such as rule generation, semantic validation, and regression checking.
+A memory store could be attached to the agent to log prior interactions with users and rewrite rules learnt over time, thus supporting continuous learning.
+The agent can decide the order in which these tools are invoked and make appropriate choices during the rewrite process.
+Finally, combining our MCTS formulation with tool-augmented agentic reasoning would also be a promising avenue for future work.
 8.4 Scope of Semantic Equivalence Tools
@@
 Therefore, a key challenge is to improve logic-based coverage.
+8.5 Enhancing NL2SQL approaches with Rewrites
+LITHE could potentially be used to enhance NL2SQL approaches to produce performant SQL queries.
+One way would be to incorporate some rewrite rules directly into NL2SQL prompts.
+Based on our empirical assessment (Figure 11(b), we would recommend the simplest Prompt 1 from the Basic Prompts in conjunction with a subset of Rules R1–R4.
+To avoid multiple prompts, one could use the LLM itself to identify the appropriate rule based on the NL query.
+While these prompts might not always provide the best rewrites, our experiments show that they are sufficient for many scenarios while incurring only small overheads as compared to using the entire prompt suite and MCTS search.
+An alternate non-invasive approach would be to simply use LITHE as a post-processor after the initial translation, but again restricted to using only a subset of the prompts/rules to keep the overheads within permissible limits.
 9 Related Work
@@
 Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
+Query rewriting in SQL engines.
+Query rewrite prior to plan enumeration has long been a feature of contemporary (non-LLM) database optimizers [9, 25].
+However, the difference in our context is that the LLM gives us the flexibility to choose from a wide variety of approaches, not restricted to the hardwired options within individual systems.
 LLMs for Database Modules.
@@
 For instance, they have been extensively used for Text-to-SQL transformations [32, 53, 4, 5, 46, 58].
+Very recently, statistical metadata was leveraged to also improve the Text-to-SQL generation process [43].
 The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [48, 40, 46].
-On the other hand, the goal of S2S rewriting is on improving the performance of an existing SQL query.
+On the other hand, the goal of SQL-to-SQL rewriting is on improving the performance of an existing SQL query.
 Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
-In recent times, LLMs have also been considered for plan-hinting [8], join-order optimization [38], index selection [49], data pipelines [22], data management [25], etc.
-An LLM-enabled multi-modal query optimizer [41], where data spans text, image, audio and video domains, has also been proposed recently.
-These approaches can be used in conjunction with the LITHE system since they address orthogonal segments of the query processing pipeline.
+In recent times, LLMs have also been considered for a variety of database modules, including plan-hinting [11], join-order optimization [47], index selection [58], data pipelines [28], data management [31], and multi-modal data optimization [50].
+These approaches can be used in conjunction with LITHE since they address orthogonal segments of the query processing pipeline.
 10 Conclusions and Future Work
@@
 Appendix A Additional Experiments
-A.1 Rewrite Quality (TPC-DS) Contd.
-Figure 10 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA for the queries run on a warm-cache environment.
-Figure 10: Execution Time Speedups in a warm-cache scenario.
-As expected, the speedups often significantly increase when compared to the cold-cache scenario.
-For instance, the speedup for Q45 increases from 700X to over $10^{4}$.
-Overall, the TSGM of LITHE increases to 17.1 (from 13.2), while that of SOTA increases to 5.7 (from 4.9).
-A.2 Rewrite Quality (Other Benchmarks)
+A.1 Rewrite Quality (Other Benchmarks)
 We also evaluated LITHE on the following set of benchmarks.
@@
 Further, the CSGM of LITHE is visibly better than SOTA.
+A.2 Rewrite Quality (Rule-based vs LLM-based approaches)
+Our SOTA baseline comprises two theoretical rule-based approaches—Learned Rewrite and LLM-R2—and two LLM-based approaches—GenRewrite and Baseline Prompt.
+The detailed breakdown of CPR achieved by these methods is presented in Table 10.
+Table 10: # CPR of Different SOTA Approaches in Comparison with LITHE
+Tools/Techniques Approach Individual # CPR >=>=1.1x
+LITHE LLM Based 26 30
+Genrewrite LLM Based 11 12
+Baseline LLM Based 10 10
+Learned Rewrite Rule Based 1 2
+LLMR2 Rule Based 0 0
+While the theoretical rule-based approaches provide formal correctness guarantees, our TPC-DS experiments revealed that the LLM-based approaches consistently outperformed them.
+Among the 13 CPRs (i.e., query rewrites with cost speedup $\geq$ 1.5X), only one was produced by a theoretical rule-based approach.
+Moreover, even when we decreased this threshold to as low as $\geq$ 1.1X, only one more rewrite became a CPR!
+LLM-based approaches, on the other hand, seem more suited for query rewrites, resulting in 12 (GenRewrite) and 30 (LITHE) rewrites with a cost speedup $\geq$ 1.1X.
+This is likely a consequence of the LLM having a “global” vision of the entire query, giving it more leeway in considering rewrites in the query-space.
+Whereas, theoretical rule-based approaches operate mainly at the operator level within the plan tree, thus being confined to only local optimizations.
+A.3 Rewrite quality (Coverage of Basic Prompts)
+Figure 11(a) shows the CPR distribution across the Basic Prompts on the Microbenchmark.
+Among the 6 rewritten queries, 3 are unique to a single prompt, while the other 3 are produced by multiple prompts.
+At first glance, one might think that Prompt 3 could be removed from our prompt suite since it does not produce any unique rewrite.
+However, this is an artifact of the micro-benchmark – on the corresponding picture for the full TPC-DS benchmark (Figure 11(b)), Prompt 3 does contribute unique rewrites.
+(a) Micro-benchmark
+(b) TPC-DS
+Figure 11: CPR distribution of Basic Prompts
 Appendix B Overheads Reduction for LITHE
@@
 In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
+As an alternative to the LLM based classifier, we also experimented with a DSPy MIPRO [38] based classifier to identify which, if any, of Rules R1–R6 is appropriate for a given query.
+However, our initial results indicated that it was weaker than our LLM-based classifier.
+While the average time taken by this approach reduced to just 10 seconds per query, the CPRs are reduced to 14, and the CSGM and TSGM fall down to 3.3 and 4.5 respectively.
+But we hasten to add that this outcome is likely an artifact of our small rule set – looking to the future, where there may be a larger corpus, a data-driven selection method like MIPRO may come to the fore.
 Pruning in MCTS
@@
     To cover predicate boundary conditions, we inject query-specific synthetic tuples into the sample—similar to the approach used by the XData mutant-killing tool [42]—in order to reduce the likelihood of non-equivalent queries producing the same results.
-Appendix D Detailed MCTS Algorithm
-Algorithm 2 Token-augmented Rewrite
-1:$root$,          # Start State;
-2:$k$,              # Maximum number of child expansion
-3:$\theta$,              # Probability threshold for node expansion
-4:$iter_{max}$,     # Maximum number of iterations
-5:$Potential$, $visits$, $V$leftarrow empty Map
-6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
-7:    $visits[root]$leftarrow $visits[root]$ + 1
-8:    $n_{cur}$leftarrow $root$
-9:    # Stage 1: Selection
-10:    while $len(n_{cur}.children)>0$ do
-11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
-12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
-13:    # Stage 2: Expansion
-14:    $expand\leftarrow$ True
-15:    while $expand$ do
-16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
-17:        if $tokens_{next}[0]=$‘;’ then
-18:           $expand\leftarrow$ False
-19:        else if $P_{next}\leq\theta$ then
-20:           for $token\in tokens_{next}$ do
-21:               $state\leftarrow n_{cur}.state\cdot token$
-22:               $n_{new}\leftarrow$ new $Node$ with State $state$
-23:               Append $n_{new}$ to $n_{cur}.children$
-24:           $expand\leftarrow$ False
-25:        else
-26:           $state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$ First Token
-27:           $n_{new}\leftarrow$ new $Node$ with State $state$
-28:           Append $n_{new}$ to $n_{cur}.children$
-29:           $n_{cur}\leftarrow n_{new}$
-30:    # Stage 3: Simulation
-31:    $query\leftarrow$ GreedyExpand($n_{cur}$)
-32:    # Compute the potential of query (returns 0 for invalid query)
-33:    $v\leftarrow$ ComputePotential($query$)
-34:    $Potential[query]\leftarrow v$
-35:    # Stage 4: Backpropagation
-36:    $n\leftarrow n_{cur}$
-37:    while $n\neq$ Null do
-38:        if $v>V[n]$ then
-39:           $V[n]\leftarrow v$
-40:        $n\leftarrow\text{Parent}(n)$
-41:# Return valid rewrite with maximum $Potential>1$
-42:if $\exists q\in Potential\mid Potential[q]>1$ then
-43:    return $q$ having maximum value of $Potential[q]$
-44:else
-45:    return the original query
-Appendix E Examples used in Prompts for Rules 1–6
+Appendix D Examples used in Prompts for Rules 1–6
 R1: Use CTEs (Common Table Expressions) to avoid repeated computation.